### v2

# On Ball’s conjectured Santaló type inequality

Abstract
We prove that if $K$ is a symmetric and isotropic convex body in $\mathbb{R}^{n}$, then
$\int_{K}\langle x,u\rangle^{2}\,dx\int_{K^{\circ}}\langle x,u\rangle^{2}\,dx\leq\left(\int_{B_{2}^{n}}\langle x,u\rangle^{2}\,dx\right)^{2},\qquad\forall u\in\mathbb{R}^{n},$
with equality for some $u\neq o$, if and only if $K$ is a Euclidean ball. This confirms a conjecture by Keith Ball (1986), stating that for any symmetric convex body $K$ in $\mathbb{R}^{n}$, it holds
$\int_{K}\int_{K^{\circ}}\langle x,y\rangle^{2}\,dx\,dy\leq\int_{B_{2}^{n}}\int_{B_{2}^{n}}\langle x,y\rangle^{2}\,dx\,dy,$
with equality if and only if $K$ is an ellipsoid.
Fortunately, our method for proving Ball’s conjectured inequality admits a quantitative stability refinement, which in turn yields an asymptotically optimal stability version of the Blaschke-Santaló inequality for origin symmetric convex bodies in terms of the symmetric difference metric. This resolves another well known open problem.
2020 Mathematics Subject Classification. Primary: 52A20; Secondary: 52A38, 52A40.Keywords. Blaschke-Santaló inequality, Ball’s conjecture, duality
1 Introduction
For each $n\in\mathbb{N}$, we fix an inner product $\langle\cdot,\cdot\rangle$ and an orthonormal basis $\{e_{1},\dots,e_{n}\}$ in $\mathbb{R}^{n}$. The origin will be denoted by $o$. For $x\in\mathbb{R}^{n}$, set $x_{j}:=\langle x,e_{j}\rangle$, $j=1,\dots,n$, and $\|x\|:=\sqrt{\langle x,x\rangle}$. Set also $B_{2}^{n}:=\{x\in\mathbb{R}^{n}:\|x\|\leq 1\}$ and $S^{n-1}:=\{x\in\mathbb{R}^{n}:\|x\|=1\}$ to be the Euclidean unit ball and the Euclidean unit sphere, respectively, with respect to the inner product $\langle\cdot,\cdot\rangle$. and denote $\omega_{n}=|B^{n}_{2}|$. Set $\mathbb{R}^{n}_{+}$ to be the first orthant, that is, $\mathbb{R}^{n}_{+}=\{x\in\mathbb{R}^{n}:x_{i}\geq 0,\ i=1,\dots,n\}$. A convex body $K$ will be a compact convex set with nonempty interior. In addition, $K$ will be called symmetric, if K=KK=-K. Finally, $K$ will be called unconditional if for any $x=(x_{1},\dots,x_{n})\in K$ and for any choice $\varepsilon_{1},\dots,\varepsilon_{n}\in\{\pm 1\}$, the point $(\varepsilon_{1}x_{1},\dots,\varepsilon_{n}x_{n})$ is also contained in $K$.
For a symmetric convex body $K$ in $\mathbb{R}^{n}$, its polar body $K^{\circ}$ is defined by
$K^{\circ}=\{x\in\mathbb{R}^{n}:\langle x,y\rangle\leq 1\;\forall y\in K\}$
and satisfies
$(TK)^{\circ}=T^{-t}K^{\circ},\qquad\forall T\in GL(n)$ (1)
One of the cornerstones of convex geometry is the classical Blaschke-Santaló inequality, which we only state here in the symmetric case.
Theorem A (Blaschke-Santaló inequality).
If $K$ is a symmetric convex body in $\mathbb{R}^{n}$, then
$|K|\cdot|K^{\circ}|\leq|B_{2}^{n}|^{2},$ (2)
with equality if and only if $K$ is an ellipsoid.
The inequality for all $n\geq 2$ is due to Santaló [31]. By now, there are several proofs of the Blaschke-Santaló inequality and some extensions of it, see e.g. [1, Theorem 7.3], [4], [5], [12], [21], [22], [25], [27], [28] and others.
Keith Ball [5], [6] has conjectured the following Santaló type inequality.
Conjecture 1.1 (Ball, 1986).
If $K$ is a symmetric convex body in $\mathbb{R}^{n}$, then
$\int_{K}\int_{K^{\circ}}\langle x,y\rangle^{2}\,dx\,dy\leq\int_{B_{2}^{n}}\int_{B^{n}_{2}}\langle x,y\rangle^{2}\,dx\,dy.$ (3)
He proved that (3) is stronger than (2), in the sense that if (3) was known to be true, then (2) would follow by the latter in a few lines.
Ball proved Conjecture 1.1 in the unconditional case. The key to his proof is the following.
Theorem B (Ball).
If $X,Y\subseteq\mathbb{R}^{n}_{+}$ are compact sets of positive volume such that $\langle x,y\rangle\leq 1$ holds for all $x\in X$ and $y\in Y$, then for any $j=1,\dots,n$, it holds
$\int_{X}x_{j}^{2}\,dx\int_{Y}x_{j}^{2}\,dx\leq\left(\int_{\mathbb{R}^{n}_{+}\cap B_{2}^{n}}x_{j}^{2}\,dx\right)^{2}=\frac{1}{2^{2n}}\left(\int_{B_{2}^{n}}x_{j}^{2}\,dx\right)^{2},$ (4)
with equality if and only if there exists an unconditional ellipsoid $E$ such that $X=\mathbb{R}^{n}_{+}\cap E$ and $Y=\mathbb{R}^{n}_{+}\cap E^{\circ}$.
To be more precise, (4) was originally established (as part of the proof of the unconditional case of Conjecture 1.1) in [5] and [6] for $X=C\cap\mathbb{R}^{n}_{+}$ and $Y=C^{\circ}\cap\mathbb{R}^{n}_{+}$, where $C$ is an unconditional convex body. However, the extension to the more general setting is straightforward (see e.g. Kalantzopoulos, Saroglou [18, Proposition 2.1] and the comments after it). We recall the argument for Theorem B, together with the characterization of the equality case in Section 2.
Our main goal is to confirm Ball’s conjecture (Conjecture 1.1) in full generality.
Theorem 1.2.
If $K$ is a symmetric convex body in $\mathbb{R}^{n}$, then
$\int_{K}\int_{K^{\circ}}\langle x,y\rangle^{2}\,dx\,dy\leq\int_{B_{2}^{n}}\int_{B^{n}_{2}}\langle x,y\rangle^{2}\,dx\,dy,$ (5)
with equality if and only if $K$ is an origin symmetric ellipsoid.
In fact, we prove a slightly stronger result in Theorem 1.3. In this paper, we say that a symmetric convex body $L\subseteq\mathbb{R}^{n}$ is isotropic (see Milman, Pajor [29] or Artstein-Avidan, Giannopoulos, Milman [2, 3]) if
$\int_{L}\langle x,u\rangle^{2}\,dx=\frac{\|u\|^{2}}{n}\int_{L}\|x\|^{2}\,dx\qquad\forall u\in\mathbb{R}^{n}.$ (6)
Note that unlike in the usual definition of isotropicity, we do not require any normalization on $L$.
Theorem 1.3.
Let $K\subseteq\mathbb{R}^{n}$ be a symmetric convex body, and let $u\in\mathbb{R}^{n}\setminus\{o\}$. If $K^{\circ}$ (or equivalently, if $K$) is isotropic, then
$\int_{K}\langle x,u\rangle^{2}\,dx\int_{K^{\circ}}\langle x,u\rangle^{2}\,dx\leq\left(\int_{B_{2}^{n}}\langle x,u\rangle^{2}\,dx\right)^{2},$ (7)
with equality if and only if $K$ is a Euclidean ball centered at the origin.
Our proof of Theorem 1.3 is inspired by Lehec’s proof [21] of a functional version of the classical Blaschke-Santaló inequality, originally established in Fradelizi-Meyer [12]. As an immediate corollary of Theorem 1.3, one has
$\int_{K}\|x\|^{2}\,dx\cdot\int_{K^{\circ}}\|x\|^{2}\,dx\leq\left(\int_{B_{2}^{n}}\|x\|^{2}\,dx\right)^{2},$ (8)
where $K$ (or $K^{\circ}$) is a symmetric isotropic convex body and equality holds if and only if $K$ is a Euclidean ball centered at the origin.
We mention that, as observed in [16, Proof of Proposition 3.1] (see also [18]), one can deduce a functional version of Theorem 1.2 by Theorem 1.2 itself. This is done by adapting the proof of [12, Proposition 3] to the setting of Conjecture 1.1.
Corollary 1.4.
Let $\rho:\mathbb{R}\to\mathbb{R}_{+}$ and $f,g:\mathbb{R}^{n}\to\mathbb{R}_{+}$ be even measurable functions satisfying $f(x)g(y)\leq\rho^{2}(\langle x,y\rangle)$, for all $x,y\in\mathbb{R}^{n}$ such that $\langle x,y\rangle>0$. Then,
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\langle x,y\rangle^{2}f(x)g(y)\,dx\,dy\leq n^{-1}\left(\int_{\mathbb{R}^{n}}\|u\|_{2}^{2}\rho\left(\|u\|_{2}^{2}\right)\,du\right)^{2}.$
Equality holds if and only if there exists a continuous function $\tilde{\rho}$$\mathbb{R}_{+}\to\mathbb{R}_{+}$, satisfying the following.
a. $\rho=\tilde{\rho}$ a.e., $\sqrt{\tilde{\rho}(s)\tilde{\rho}(t)}\leq\tilde{\rho}(\sqrt{st})$ for every $s,t\geq 0$ and if $n\geq 2$, then $\tilde{\rho}(0)>0$ or $\tilde{\rho}|_{\mathbb{R}_{+}}\equiv 0$.
b. For some positive definite $n\times n$ matrix $T$ and for some $d>0$, one has
$f(x)=d\tilde{\rho}(|Tx|^{2})\hbox{ and }g(x)=\frac{1}{d}\tilde{\rho}(|T^{-1}x|^{2})\quad a.e.$
Fusco, Maggi, Pratelli [14] proved an optimal stability version of the Isoperimetric inequality in terms of the symmetric difference metric, which result was extended to the Anisotropic Isoperimetric inequality (or equivalently, to the Brunn-Minkowski inequality) by Figalli, Maggi, Pratelli [13]. For symmetric convex bodies $K,C\subseteq\mathbb{R}^{n}$, their homothetic distance, considered, for example, by Figalli, Maggi, Pratelli [13] and Fusco, Maggi, Pratelli [14], is
$A(K,C)=|\alpha K\Delta\beta C|\mbox{ \ where }\alpha=|K|^{-\frac{1}{n}}\mbox{ and }\beta=|C|^{-\frac{1}{n}}.$
Here $|\alpha K|=|\beta C|=1$. It is easy to see that the Banach-Mazur distance
$d(K,B^{n}_{2})=\min\{\log\lambda:\,\lambda\geq 1\mbox{ and }\exists\Phi\in{\rm GL}(n),\;\Phi K\subseteq B^{n}_{2}\subseteq\lambda\Phi K\}$
of $K$ and $B^{n}_{2}$ satisfies
$d(K,B^{n}_{2})\leq C_{n}A(K,E)^{\frac{2}{n+1}}$ (9)
for any centered ellipsoid $E$ where $C_{n}>0$ depends on $n$. We are ready to state the stability version Theorem 1.5 of Theorem 1.2.
Theorem 1.5.
If $K\subseteq\mathbb{R}^{n}$ is a symmetric convex body, then there exists an $o$-symmetric ellipsoid $E\subseteq\mathbb{R}^{n}$ such that
$\int_{K}\int_{K^{\circ}}\langle x,y\rangle^{2}\,dydx\leq\left(1-\theta_{n}A(K,E)^{2}\right)\int_{B^{n}_{2}}\int_{B^{n}_{2}}\langle x,y\rangle^{2}\,dydx$ (10)
where $\theta_{n}>0$ is an explicit constant depending only on $n$.
Theorem 1.5 implies the stability of various additional linear invariant fundamental inequalities in terms of the symmetric difference metrics. Stability versions of the Blaschke-Santaló inequality and the Affine Isoperimetric inequality (cf. (12)) have been known in terms of the ”easier to handle” Banach-Mazur-distance (see Ball, Böröczky [7], Böröczky [9] and Ivaki [17]), but this is the first instance, following the footsteps of Fusco, Maggi, Pratelli [14] and Figalli, Maggi, Pratelli [13], to provide a stability version symmetric difference metrics, which is the most natural distance in this respect.
Theorem 1.6.
If $K\subseteq\mathbb{R}^{n}$ is a symmetric convex body, then there exists an $o$-symmetric ellipsoid $E\subseteq\mathbb{R}^{n}$ such that
$|K|\cdot|{\color[rgb]{0,0,0}K^{\circ}}|\leq\left(1-\theta_{n}A(K,E)^{2}\right)\omega_{n}^{2},$ (11)
where $\theta_{n}>0$ is an explicit constant depending only on $n$.
Here the stability estimate with respect to the Banach-Mazur distance after Theorem 1.6 is an improvement on the best one previously known provided by Ball, Böröczky [7].
It is well-known (cf. Lutwak [26]) that the Blaschke-Santaló inequality (2) is equivalent to the Affine Isoperimetric inequality (12). For a symmetric convex body $K\subseteq\mathbb{R}^{n}$, the Gaussian curvature $\kappa_{K}(x)\geq 0$ exists at $\mathcal{H}^{n-1}$ a.e. $x\in\partial K$ by Alexandrov’s theorem, and the affine surface area is defined as
$\Omega(K)=$ $\int_{\partial K}\kappa_{K}(x)^{\frac{1}{n+1}}\,d\mathcal{H}^{n-1}(x)$
According to Ludwig [24] (crowning the developments due to Lutwak [25], Schütt, Werner [33] and Dolzmann, Hug [10]) the affine surface area is equi-linear invariant ($\Omega(\Phi K)=\Omega(K)$ for $\Phi\in{\rm GL}(n)$ with $\det\Phi=\pm 1$), it is upper semi-continuous and it satisfies the Affine Isoperimetric inequality saying that
$\Omega(K)\leq$ $n\omega_{n}^{\frac{2}{n+1}}|K|^{\frac{n-1}{n+1}}$ (12)
with equality if and only if $K$ is an ellipsoid. Now, Theorem 1.6 implies the first close to be optimal stability version of the Affine Isoperimetric inequality in terms of the symmetric volume difference.
Corollary 1.7.
If $K\subseteq\mathbb{R}^{n}$ is a symmetric convex body, then there exists an $o$-symmetric ellipsoid $E\subseteq\mathbb{R}^{n}$ such that
$\Omega(K)\leq$ $\left(1-\theta_{n}A(K,E)^{2}\right)n\omega_{n}^{\frac{2}{n+1}}|K|^{\frac{n-1}{n+1}}$ (13)
where $\theta_{n}>0$ is an explicit constant depending only on $n$.
Remark. Concerning the Banach-Mazur distance, it follows from (9) that in Theorems 1.5 and 1.6 and in Corollary 1.7, the term $\theta_{n}A(K,E)^{2}$ may be replaced by $\tilde{\theta}_{n}d(K,B_{2}^{n})^{n+1}$ where $\tilde{\theta}_{n}>0$ depends on $n$.
Here the stability estimates with respect to the Banach-Mazur distance after Theorem 1.7 are improvements on the best ones previously known, proved by Ball, Böröczky [7].
This paper can be naturally divided into two parts. The first part (Sections 2, 3, 4) is devoted to the proof of Ball’s conjecture (Theorems 1.3 and 1.2). The second part (in which some of the arguments of the first part are repeated), consisting of Sections 5, 6, 7, proves the stability results of the paper.
2 Proof of Theorem B
The proof of Theorem B will be a consequence of the classical Prékopa-Leindler inequality, where we only quote the equality case in the case we need. This argument will have important role also in proving the stability version Theorem 1.6 of Keith Ball’s conjectured inequality.
Theorem 2.1 (Prékopa-Leindler-Dubuc [30], [23], [11]).
If $h,f,g:\,\mathbb{R}^{n}\to[0,\infty)$ are integrable functions on $\mathbb{R}^{n}$ with positive integral satisfying $h((x+y)/2))\geq\sqrt{f(x)g(y)}$ for any $x,y\in\mathbb{R}^{n}$, then
$\left(\int_{\mathbb{R}^{n}}h\right)^{2}\geq\left(\int_{\mathbb{R}^{n}}f\right)\cdot\left(\int_{\mathbb{R}^{n}}g\right).$ (14)
If, in addition, $h$ is log-concave, and equality holds in (14), then there exists $w\in\mathbb{R}^{n}$ and $a>0$, such that f(x)=ah(xw)f(x)=ah(x-w) and $g(x)=a^{-1}h(x+w)$ for a.e. $x\in\mathbb{R}^{n}$.
For $X,Y\subseteq\mathbb{R}^{n}_{+}$, we consider the “coordinatewise geometric mean”
$X^{\frac{1}{2}}\cdot Y^{\frac{1}{2}}=\{(\sqrt{x_{1}y_{1}},\dots,\sqrt{x_{n}y_{n}})\in\mathbb{R}^{n}_{+}:(x_{1},\dots,x_{n})\in X,\ (y_{1},\dots,y_{n})\in Y\}.$
We observe that if $\langle x,y\rangle\leq 1$ for any $x\in X$ and $y\in Y$, then
$X^{\frac{1}{2}}\cdot Y^{\frac{1}{2}}\subseteq B^{n}_{2}.$ (15)
Proof of Theorem B..
We may assume that $i=1$. Since $X\cap{\rm int}\,\mathbb{R}^{n}_{+}\neq\emptyset$ and $Y\cap{\rm int}\,\mathbb{R}^{n}_{+}\neq\emptyset$, the condition $\langle x,y\rangle\leq 1$ for any $x\in X$ and $y\in Y$ yields that both $X$ and $Y$ are bounded. We consider the integrable functions $h,f,g:\,\mathbb{R}^{n}\to[0,\infty)$ defined as
$f(t_{1},\dots,t_{n})=$ $e^{2t_{1}}\mathbf{1}_{X}(e^{t_{1}},\dots,e^{t_{n}})\cdot e^{t_{1}+\dots+t_{n}},$ (16)
$g(t_{1},\dots,t_{n})=$ $e^{2t_{1}}\mathbf{1}_{Y}(e^{t_{1}},\dots,e^{t_{n}})\cdot e^{t_{1}+\dots+t_{n}},$ (17)
$h(t_{1},\dots,t_{n})=$ $e^{2t_{1}}\mathbf{1}_{B^{n}_{2}}(e^{t_{1}},\dots,e^{t_{n}})\cdot e^{t_{1}+\dots+t_{n}},$ (18)
and hence the substitution $(t_{1},\dots,t_{n})\mapsto(e^{t_{1}},\dots,e^{t_{n}})$ shows that these functions satisfy
$\int_{\mathbb{R}^{n}}f(x)\,dx=\int_{X}x_{1}^{2}\,dx,\qquad\int_{\mathbb{R}^{n}}g(x)\,dx=\int_{Y}x_{1}^{2}\,dx,\qquad\int_{\mathbb{R}^{n}}h(x)\,dx=\int_{B^{n}_{2}}x_{1}^{2}\,dx.$ (19)
Moreover, (15) and the condition $\langle x,y\rangle\leq 1$ for any $x\in X$ and $y\in Y$ ensure that
$h\left((x+y)/2)\right)\geq\sqrt{f(x)g(y)},$
for any $x,y\in\mathbb{R}^{n}$. Therefore, the Prékopa-Leindler inequality (14) yields (4).
Assume now that equality holds in (4). Thus, equality holds in the Prékopa-Leindler inequality (14) for $f,g,h$. We deduce from Theorem 2.1 that there exist $a>0$ and $w=(w_{1},\dots,w_{n})\in\mathbb{R}^{n}$ such that for $b_{j}=e^{w_{j}}$, $j=1,\dots,n$, we have
$e^{2t_{1}}\mathbf{1}_{X}(e^{t_{1}},\dots,e^{t_{n}})\cdot e^{t_{1}+\dots+t_{n}}=$ $a\cdot\frac{e^{2t_{1}}}{b_{1}^{2}}\cdot\mathbf{1}_{B^{n}_{2}}\left(\frac{e^{t_{1}}}{b_{1}},\dots,\frac{e^{t_{n}}}{b_{n}}\right)\cdot\frac{e^{t_{1}+\dots+t_{n}}}{b_{1}\dots b_{n}},$
$e^{2t_{1}}\mathbf{1}_{Y}(e^{t_{1}},\dots,e^{t_{n}})\cdot e^{t_{1}+\dots+t_{n}}=$ $b_{1}^{2}\cdot\frac{e^{2t_{1}}}{a}\cdot\mathbf{1}_{B^{n}_{2}}\left(b_{1}e^{t_{1}},\dots,b_{n}e^{t_{n}}\right)\cdot e^{t_{1}+\dots+t_{n}}\cdot b_{1}\dots b_{n}.$
As the value of an indicator function is 0 or $1$, and $X$ and $Y$ are compact, it follows that
$X=\mathbb{R}^{n}_{+}\cap\Phi B^{n}_{2}\qquad\textnormal{and}\qquad Y=\mathbb{R}^{n}_{+}\cap\Phi^{-1}B^{n}_{2},$
for the diagonal transformation $\Phi$ with eigenvalues $b_{1},\dots,b_{n}$. This proves Theorem B. ∎
3 Yao-Yao partitions
Let $E$ be an affine space of finite dimension. For the purposes of this paper, a partition of $E$ will be a finite family of non overlapping convex cones whose union is the whole $E$. A special family of partitions, called Yao-Yao partitions, was invented by Yao and Yao [34]. A Yao-Yao partition $\cal P$} (together with its center) is defined inductively as follows.
Definition 3.1.
Let $E$ be an affine subspace of a finite dimensional vector space $V$. If $\dim E=0$, i.e. $E=\{x\}$ for some element $x\in V$, the family ${\cal P}=\{x\}$ is defined to be a Yao-Yao partition of $E$ and $x$ is called the center of $\cal P$}. When $\dim E=n\geq 1$, we say that $\cal P$} is a Yao-Yao partition of $E$ with center $x\in E$, if there exists a hyperplane $F$ of $E$, a vector $v\in V$ which is not parallel to $F$ and two Yao-Yao partitions $\cal P$$_{+}$ and $\cal P$$_{-}$ of $F$ with the same center $x\in F$, such that
${\cal P}=\{A+\mathbb{R}_{+}v:A\in{\cal P}_{+}\}\cup\{A+\mathbb{R}_{-}v:A\in{\cal P}_{-}\}.$
In this case, we say that the Yao-Yao partition $\cal P$} is based on the hyperplane $F$ of $E$, and $v$ is an axis of $\cal P$}.
Let us collect the basic properties of a Yao-Yao partition $\cal P$} of $\mathbb{R}^{n}$, that will be needed later (see [20]).
  • The center of $\cal P$} is well defined.
  • $\cal P$} consists of $2^{n}$ elements of the form $x+\textnormal{pos}\,\{v_{1},\dots,v_{n}\}$, where $x$ is the center of $\cal P$} and $v_{1},\dots,v_{n}\in\mathbb{R}^{n}$ are linearly independent vectors.
  • If the center of $\cal P$} is the origin, then the family ${\cal P}^{*}:=\{A^{*}:A\in{\cal P}\}$ is also a partition of $\mathbb{R}^{n}$ (see [21]). Here,
    $A^{*}:=\{y\in\mathbb{R}^{n}:\langle x,y\rangle\geq 0,\ \forall x\in A\}$
    is the dual cone of $A$.
  • If $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a non-singular affine map, then the family $\{TA:A\in{\cal P}\}$ is also a Yao-Yao partition of $\mathbb{R}^{n}$.
  • If $H^{+}\subseteq E$ is a half-space containing the center of $\cal P$}, then there exists a $A\in{\cal P}$ such that
    $A\subseteq{\color[rgb]{0,0,0}H^{+}}.$ (20)
Let $\mu$ be a finite Borel measure on $\mathbb{R}^{n}$. We say that a Yao-Yao partition P is a Yao-Yao equipartition of $\mu$, if for all $A\in{\cal P}$, $\mu(A)=2^{-n}\mu(\mathbb{R}^{n})$. Under some mild assumptions, the existence of a Yao-Yao equipartition for $\mu$ was established by Yao and Yao [34] and later by Lehec [20] (under weaker assumptions on $\mu$). Lehec’s construction appears to provide more information than the original construction by Yao and Yao. For example, let $\lambda_{1},\dots,\lambda_{n}:\mathbb{R}^{n}\to\mathbb{R}$ be affine forms, such that the map
$E\in z\mapsto(\lambda_{1}(z),\dots,\lambda_{n}(z))\in\mathbb{R}^{n}$
is one to one. Following [20, Definition 6], we say (inductively) that a Yao-Yao partition $\cal P$} based on a hyperplane $F\subseteq\mathbb{R}^{n}$ is adapted to the coordinate system $(\lambda_{1},\dots,\lambda_{n})$ if $F=\{y\in E:\lambda_{1}(y)=\lambda_{1}(x)\}$ and $\cal P$$_{+}$, $\cal P$$_{-}$ are both adapted to $(\lambda_{2}|_{F},\dots,\lambda_{n}|_{F})$. According to Lehec’s [20, Proposition 14], if $\mu$ is a finite Borel measure on $\mathbb{R}^{n}$ such that $\mu(H)=0$ for all hyperplanes $H\subseteq\mathbb{R}^{n}$ and $\textnormal{supp}\,\mu=\mathbb{R}^{n}$, then there exists a Yao-Yao equipartition of $\mu$ that is adapted to $(\lambda_{1},\dots,\lambda_{n})$, and two such Yao-Yao partitions have the same center. In particular, the following crucial ingredient for the proof of Theorem 1.3 can be deduced from the results in [20].
Theorem C (Lehec).
Let $(\lambda_{1},\dots,\lambda_{n})$ be a system of coordinates and $\mu$ an even finite Borel measure on $\mathbb{R}^{n}$ such that $\mu(\mathbb{R}^{n})>0$ and $\mu(H)=0$ holds for all hyperplanes $H\subseteq\mathbb{R}^{n}$. Then there exists a Yao-Yao equipartition $\cal{P}$ of $\mu$ adapted to $(\lambda_{1},\dots,\lambda_{n})$ having its center at the origin. In particular, for any $u\in S^{n-1}$, there exists a Yao-Yao equipartition $\cal P$} of $\mu$, which is based on $u^{\perp}$ and has its center at the origin.
Proof.
For an even and integrable function $f:\mathbb{R}^{n}\to(0,\infty)$ and for a positive integer $k$, let $\mu_{k}:=\mu+(1/k)f\,dx$, and hence $\mu_{k}$ is an even finite Borel measure that assigns zero mass to any hyperplane and $\textnormal{supp}\,\mu_{k}=\mathbb{R}^{n}$. According to [20, Proposition 14], there exists a unique Yao-Yao equipartition ${\cal P}_{k}$ of $\mu_{k}$ which is adapted $(\lambda_{1},\dots,\lambda_{n})$. Since $\mu_{k}$ is also even, uniqueness of the center ensures that for each $k$, the center of ${\cal P}_{k}$ is the origin (cf. [20, Lemma 15]). But then, approximation ([20, Lemma 12]) shows that $\mu$ also has a Yao-Yao equipartition $\cal P$} with center $o$, adapted to $(\lambda_{1},\dots,\lambda_{n})$. In particular, taking $\lambda_{1}(x)=\langle u,x\rangle$, $\cal P$} is based on the hyperplane
$\{y\in\mathbb{R}^{n}:\,\lambda_{1}(y)=\lambda_{1}(o)\}=u^{\perp},$
which is exactly what we wanted to show. ∎
4 Proof of Theorem 1.2 and Theorem 1.3
Proof of Theorem 1.3.
To prove (7), we may clearly assume that $u=e_{1}$. Define the measure $\mu:={\bf 1}_{K}x_{1}^{2}dx$. By Theorem C, there are two Yao-Yao partitions $\cal P$$_{+}$, $\cal P$$_{-}$ of $e_{1}^{\perp}$ and a vector $v\in S^{n-1}$ with $\langle v,e_{1}\rangle>0$, such that the partition
${\cal P}:=\{A^{\prime}+\mathbb{R}_{+}v:A^{\prime}\in{\cal P}_{+}\}\cup\{A^{\prime}+\mathbb{R}_{-}v:A^{\prime}\in{\cal P}_{-}\}$
is a Yao-Yao equipartition of $\mu$ with center $o$. Let $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be the linear map, such that $T(v)=\langle e_{1},v\rangle e_{1}$ and $Tz=z$ for $z\in e_{1}^{\perp}$, and hence $\det T=1$. Then, for an arbitrary $x\in\mathbb{R}^{n}$ we have
$\langle T^{-1}x,e_{1}\rangle=x_{1}\langle T^{-1}e_{1},e_{1}\rangle+\langle T^{-1}(x-x_{1}e_{1}),e_{1}\rangle=x_{1}\langle v/\langle e_{1},v\rangle,e_{1}\rangle=x_{1}.$ (21)
Set $\widetilde{K}:=TK$. Using (21) for any Borel set $\Omega$ in $\mathbb{R}^{n}$, we get
$\int_{(T\Omega)\cap\widetilde{K}}x_{1}^{2}\,dx=\int_{(T\Omega)\cap\widetilde{K}}\langle T^{-1}x,e_{1}\rangle^{2}\,dx=\int_{T^{-1}((T\Omega)\cap\widetilde{K})}\langle x,e_{1}\rangle^{2}\,dx=\int_{\Omega\cap K}x_{1}^{2}\,dx.$
In particular, we conclude that
$\int_{\widetilde{K}}x_{1}^{2}\,dx=\int_{K}x_{1}^{2}\,dx$ (22)
and if ${\cal P}^{\prime}:=\{TA:A\in{\cal P}\}$, then $\cal P$$^{\prime}$ is a Yao-Yao equipartition for $\tilde{\mu}:={\bf 1}_{\widetilde{K}}x_{1}^{2}dx$ based on $e_{1}^{\perp}$ as
$\int_{B\cap\widetilde{K}}x_{1}^{2}\,dx=\frac{1}{2^{n}}\int_{\widetilde{K}}x_{1}^{2}\,dx,\qquad\forall B\in{\cal P}^{\prime}.$ (23)
On the other hand, we have
$\int_{{\widetilde{K}}^{\circ}}x_{1}^{2}\,dx=\int_{T^{-t}K^{\circ}}x_{1}^{2}\,dx=\int_{K^{\circ}}\langle T^{-t}x,e_{1}\rangle^{2}\,dx=\int_{K^{\circ}}\langle x,T^{-1}e_{1}\rangle^{2}\,dx=\frac{1}{\langle e_{1},v\rangle^{2}}\int_{K^{\circ}}\langle x,v\rangle^{2}\,dx.$ (24)
Hence, $|\langle e_{1},v\rangle|\leq 1$ and the isotropicity of $K^{\circ}$ (cf. (6)) imply
$\int_{{\widetilde{K}}^{\circ}}x_{1}^{2}\,dx\geq\int_{K^{\circ}}\langle x,v\rangle^{2}\,dx=\int_{K^{\circ}}x_{1}^{2}\,dx.$ (25)
Now, (22) and (25) yield
$\int_{K}x_{1}^{2}\,dx\int_{K^{\circ}}x_{1}^{2}\,dx\leq\int_{\widetilde{K}}x_{1}^{2}\,dx\int_{{\widetilde{K}}^{\circ}}x_{1}^{2}\,dx.$ (26)
Clearly, equality holds in (26) if and only if $v=e_{1}$. It remains to prove that (7) holds with $u=e_{1}$ and with $\widetilde{K}$} in the place of $K$. To this end, observe that if $B\in{\cal P}^{\prime}$, then there exist $C\subseteq e_{1}^{\perp}$ and $\varepsilon\in\{-1,1\}$, such that $B=C+\varepsilon\mathbb{R}_{+}e_{1}$, where $C$ is of the form $\textnormal{pos}\,\{v_{1},\dots,v_{n-1}\}$, for some linearly independent vectors $v_{1},\dots,v_{n-1}\in e_{1}^{\perp}$. We wish to show that
$\int_{B\cap{\widetilde{K}}}x_{1}^{2}\,dx\int_{B^{*}\cap{\widetilde{K}}^{\circ}}x_{1}^{2}\,dx\leq\frac{1}{2^{2n}}\left(\int_{B_{2}^{n}}x_{1}^{2}\,dx\right)^{2},$ (27)
with equality if and only if for for this particular $B\in{\cal P}^{\prime}$, there exists a $o$-symmetric ellipsoid $E_{B}$, which is symmetric with respect to $e_{1}^{\perp}$ and satisfies $B\cap E_{B}={\color[rgb]{0,0,0}B\cap{\widetilde{K}}}$.
By symmetry, we may assume that $\varepsilon=1$. Let $S:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be any linear map (which depends on $B$) with $S(e_{1})=e_{1}$, $S(C)=\mathbb{R}^{n}_{+}\cap e_{1}^{\perp}$ and $|\det S|=1$. Then, clearly,
$S(B)=\mathbb{R}^{n}_{+}=(S(B))^{*}=S^{-t}(B^{*}).$
Notice also that, for $x\in\mathbb{R}^{n}$, it holds $\langle Sx,e_{1}\rangle=x_{1}$. Consequently, (23) gives
$\int_{\mathbb{R}^{n}_{+}\cap(S{\widetilde{K}})}x_{1}^{2}\,dx$ $=$ $\int_{S(B\cap{\widetilde{K}})}x_{1}^{2}\,dx=\int_{B\cap{\widetilde{K}}}\langle Sx,e_{1}\rangle^{2}\,dx=\int_{B\cap{\widetilde{K}}}x_{1}^{2}\,dx$ (28)
$=$ $\frac{1}{2^{n}}\int_{\widetilde{K}}x_{1}^{2}\,dx.$ (29)
In addition,
$\int_{\mathbb{R}^{n}_{+}\cap(S{\widetilde{K}})^{\circ}}x_{1}^{2}\,dx$ $=$ $\int_{S^{-t}(B^{*}\cap{\widetilde{K}}^{\circ})}x_{1}^{2}\,dx=\int_{B^{*}\cap{\widetilde{K}}^{\circ}}\langle S^{-t}x,e_{1}\rangle^{2}=\int_{B^{*}\cap{\widetilde{K}}^{\circ}}\langle x,S^{-1}e_{1}\rangle^{2}\,dx$ (30)
$=$ $\int_{B^{*}\cap{\widetilde{K}}^{\circ}}x_{1}^{2}\,dx.$
Now, (27) follows immediately from (28), (30) and Theorem B. In fact, if equality holds in (27), then equality necessarily holds in (4) with $S{\widetilde{K}}$ in place of $K$. In other words, there exists an unconditional ellipsoid $E$, such that $\mathbb{R}^{n}_{+}\cap(S{\widetilde{K}})=\mathbb{R}^{n}_{+}\cap E$. Observe that the ellipsoid $E_{B}:=S^{-1}(E)$ is symmetric with respect to $e_{1}^{\perp}$ and that $B\cap E_{B}=B\cap{\widetilde{K}}$. This establishes the characterization of equality cases in (27), claimed earlier.
Recall that the family $({\cal P}^{\prime})^{*}=\{B^{*}:B\in{\cal P}^{\prime}\}$ is also a partition of $\mathbb{R}^{n}$. Therefore, (27) and (29) give
$\int_{\widetilde{K}}x_{1}^{2}\,dx\int_{{\widetilde{K}}^{\circ}}x_{1}^{2}\,dx$ $=$ $\sum_{B\in{\cal P}^{\prime}}\int_{{\widetilde{K}}\cap B}x_{1}^{2}\,dx\sum_{D\in({\cal P}^{\prime})^{*}}\int_{{\widetilde{K}}^{\circ}\cap D}x_{1}^{2}\,dx$
$=$ $2^{n}\sum_{D\in({\cal P}^{\prime})^{*}}\int_{{\widetilde{K}}\cap D^{*}}x_{1}^{2}\,dx\int_{{\widetilde{K}}^{\circ}\cap D}x_{1}^{2}\,dx$
$\leq$ $2^{n}\sum_{D\in({\cal P}^{\prime})^{*}}\frac{1}{2^{2n}}\left(\int_{B_{2}^{n}}x_{1}^{2}\,dx\right)^{2}=\left(\int_{B_{2}^{n}}x_{1}^{2}\,dx\right)^{2}.$
This completes the proof of (7).
Finally, we assume that equality holds in (7), and hence equality holds in (26) and (27) for any $B\in\mathcal{P^{\prime}}$, as well. We deduce from the equality conditions in (26) that $v=e_{1}$, and hence ${\widetilde{K}}=K$ and $\mathcal{P^{\prime}}=\mathcal{P}$. It follows from the equality conditions in (27) that for any $A\in\mathcal{P}$, there exists an $o$-symmetric ellipsoid $E_{A}\subseteq\mathbb{R}^{n}$ such that $A\cap K=A\cap E_{A}$. If $n=2$, $v=u=e_{1}$ implies $\mathcal{P}=\{\textnormal{pos}\{\varepsilon_{1}e_{1},\varepsilon_{2}e_{2}\}:\varepsilon_{i}=\pm 1\}$. By Theorem B, $\widetilde{K}$} is an unconditional ellipse, hence $K$ is an ellipse itself. The fact that $K$ is a Euclidean disk follows from the isotropicity of $K^{\circ}$.
To handle the case $n\geq 3$, we say that two cones $A,B\in\mathcal{P}$ are neighbors, if $A={\rm pos}\{A^{\prime},e_{1}\}$ and $B={\rm pos}\{B^{\prime},e_{1}\}$, where $(A^{\prime}\cap B^{\prime})\backslash{\rm pos(e_{1})}\neq\emptyset$. We claim that if $A,B\in\mathcal{P}$ are neighbors, then
$E_{A}=E_{B}.$ (32)
Take another Yao-Yao equipartition $\mathcal{Q}$ based on $e_{1}^{\perp}$, with axis $e_{1}$, that contains some $\Omega\in\mathcal{Q}$ such that $\textnormal{int}(A\cap\Omega)\neq\emptyset$ and $\textnormal{int}(B\cap\Omega)\neq\emptyset$ (the existence of such an equipartition is due to the fact that we may arbitrarily choose a hyperplane $H$ of $e_{1}^{\perp}$ such that $\mathcal{P_{+}}$ is based on $H$; recall Theorem C) . Then there exists an ellipsoid $E_{\Omega}$ such that $\Omega\cap K=\Omega\cap E_{\Omega}$. But then
$\partial(A\cap E_{A})\cap\textnormal{int}(A\cap\Omega)=\partial(\Omega\cap E_{\Omega})\cap\textnormal{int}(A\cap\Omega).$
Since the boundary of an ellipsoid is a real analytic surface, we obtain $E_{A}=E_{\Omega}$. Similarly, $E_{B}=E_{\Omega}$ proving the claim (32).
Finally, it follows via induction on the dimension that the graph $G$ whose vertices are the $2^{n-1}$ elements of $\{A\in\mathcal{P}:x_{1}\geq 0~~\forall x\in A\}$ and the edges formed by the neighboring elements is connected. Therefore, for any $A,A^{\prime}\in\mathcal{P}$, it holds $E_{A}=E_{A^{\prime}}$ and hence $K$ is an origin symmetric ellipsoid. Since $K^{\circ}$ is isotropic, the ellipsoid $K$ is a Euclidean ball. ∎
Proof of Theorem 1.2.
Let $K$ be a symmetric convex body in ${\mathbb{R}}^{n}$ and let $K^{\prime}$ be an isotropic linear image of $K$. Since the functional
$K\mapsto\int_{K}\int_{K^{\circ}}\langle x,y\rangle^{2}\,dy\,dx$
is invariant under non-singular linear maps, it suffices to prove (3) for $K^{\prime}$ instead of $K$. By the isotropicity of $K^{\prime}$, we get
$\int_{K^{\prime}}x_{i}x_{j}\,dx=0,\qquad i\neq j$
and therefore,
$\int_{K^{\prime}}\int_{(K^{\prime})^{\circ}}\langle x,y\rangle^{2}\,dy\,dx=\sum_{i=1}^{n}\int_{K^{\prime}}x_{i}^{2}\,dx\int_{(K^{\prime})^{\circ}}y_{i}^{2}\,dy.$ (33)
The proof of (3) for $K^{\prime}$ follows immediately from (33), Theorem 1.3 and then again (33) with $K^{\prime}=(K^{\prime})^{\circ}=B_{2}^{n}$. Moreover, if equality holds in (3), then equality must hold in (7) for $K^{\prime}$ in the place of $K$ and for $u=e_{1}$. It follows by the equality cases in Theorem 1.3 that $K^{\prime}$ is a Euclidean ball. That is, $K$ is an ellipsoid. ∎
5 Weak stability of Theorem 1.3
We note that if $L\subseteq\mathbb{R}^{n}$ is a symmetric convex body in isotropic position, then Kannan, Lovász, Simonovits [19] prove that there exists an $o$-symmetric ball $B\subseteq L$ such that $L\subseteq nB$. If, in addition, $|L|=\omega_{n}$, then we deduce that
[…]
Figalli, van Hintum, Tiba [15] prove the following stability version of the Prékopa-Leindler inequality (14).
Theorem 5.1 (Figalli, van Hintum, Tiba).
For some explicit constant $\theta_{n}>1$ depending on $n\geq 1$, if $h,f,g:\,\mathbb{R}^{n}\to[0,\infty)$ are integrable with positive integral and $h$ is log-concave such that $h(\frac{1}{2}\,x+\frac{1}{2}\,y)\geq\sqrt{f(x)g(y)}$ for $x,y\in\mathbb{R}^{n}$, and
$\left(\int_{\mathbb{R}^{n}}h\right)^{2}\leq(1+\varepsilon)\left(\int_{\mathbb{R}^{n}}f\right)\left(\int_{\mathbb{R}^{n}}g\right)$
for $\varepsilon\in(0,1]$, then there exists $w\in\mathbb{R}^{n}$ such that for $a=\sqrt{\int_{\mathbb{R}^{n}}g/\int_{\mathbb{R}^{n}}f}$, we have
$\int_{\mathbb{R}^{n}}|h(x)-af(x-w)|\,dx$ $\leq$ $\theta_{n}\sqrt{\varepsilon}\cdot\int_{\mathbb{R}^{n}}h$
$\int_{\mathbb{R}^{n}}|h(x)-a^{-1}g(x+w)|\,dx$ $\leq$ $\theta_{n}\sqrt{\varepsilon}\cdot\int_{\mathbb{R}^{n}}h.$
For $t>0$ and $u\in S^{n-1}$, let us consider the strip
$\Theta_{u,t}=\{(x_{1},\ldots,x_{n})\in\mathbb{R}^{n}:|\langle x,u\rangle|\leq t\}.$
Comparing a set with positive Lebesgue measure with the centered ball of the same volume shows that if $X\subseteq\mathbb{R}^{n}$ is a bounded measurable set, then
$\int_{X}\|x\|^{2}\,dx\geq\frac{n\omega_{n}^{-\frac{2}{n}}}{n+2}\cdot|X|^{\frac{n+2}{n}}.$ (35)
The key statement is the following stability version of Theorem B.
Proposition 5.2.
For measurable $X,Y\subseteq\mathbb{R}^{n}_{+}$ with $|X|>0$ and $|Y|>0$, and $i=1,\ldots,n$, if $\langle x,y\rangle\leq 1$ for any $x\in X$ and $y\in Y$, and
$\int_{X}x_{i}^{2}\;dx\cdot\int_{Y}x_{i}^{2}\;dx\geq(1-\varepsilon)\left(\int_{\mathbb{R}^{n}_{+}\cap B^{n}_{2}}x_{i}^{2}\;dx\right)^{2}$ (36)
for some $\varepsilon\in(0,\frac{1}{2})$, then there exists positive definite diagonal matrix $\Phi$ such that
$\left|(\mathbb{R}^{n}_{+}\cap B^{n}_{2})\Delta(\Phi X)\right|<$ $\theta_{n}\varepsilon^{\frac{n}{2(n+2)}}\mbox{ \ and }$ $\left|(\mathbb{R}^{n}_{+}\cap B^{n}_{2})\Delta(\Phi^{-1}Y)\right|<$ $\theta_{n}\varepsilon^{\frac{n}{2(n+2)}}$ (37)
[…]
for any $\aleph>0$ where the explicit $\theta_{n}>1$ depends on $n$.
Proof.
We may assume that $i=1$, and after rescaling $X$ and $Y$, we may assume that
$\int_{X}{\color[rgb]{0,0,0}x_{1}^{2}}\;dx=\int_{Y}x_{1}^{2}\;dx,$
and hence, using also Theorem 1.3, we have
$(1-\varepsilon)\int_{\mathbb{R}^{n}_{+}\cap B^{n}_{2}}x_{1}^{2}\;dx<\int_{X}x_{1}^{2}\;dx\leq\int_{\mathbb{R}^{n}_{+}\cap B^{n}_{2}}x_{1}^{2}\;dx.$
We consider again the functions $f,g,h$ defined in (16), (17) and (18), which satisfy (19) as well as $h(\frac{1}{2}\,x+\frac{1}{2}\,y)\geq\sqrt{f(x)g(y)}$ for any $x,y\in\mathbb{R}^{n}$. Therefore, combining (36) and the stability version Theorem 5.1 of the Prékopa-Leindler inequality yields the existence of a $w\in\mathbb{R}^{n}$ such that
$\int_{\mathbb{R}^{n}}|h(x)-f(x-w)|\,dx\leq$ $\tilde{\theta}_{n}\sqrt{\varepsilon}\cdot\int_{\mathbb{R}^{n}}h$
$\int_{\mathbb{R}^{n}}|h(x)-g(x+w)|\,dx\leq$ $\tilde{\theta}_{n}\sqrt{\varepsilon}\cdot\int_{\mathbb{R}^{n}}h$
where $\tilde{\theta}_{n}>1$ depends on $n$. For $C_{n}=\tilde{\theta}_{n}\int_{\mathbb{R}^{n}_{+}\cap B^{n}_{2}}x_{1}^{2}\;dx$ and $b_{j}=e^{w_{j}}$, $j=1,\ldots,n$, we deduce that
$\int_{(\mathbb{R}^{n}_{+}\cap B^{n}_{2})\Delta\Phi X}x_{1}^{2}\;dx\leq$ $C_{n}\sqrt{\varepsilon}$
$\int_{(\mathbb{R}^{n}_{+}\cap B^{n}_{2})\Delta\Phi^{-1}Y}x_{1}^{2}\;dx\leq$ $C_{n}\sqrt{\varepsilon}$
where $\Phi$ is the diagonal transformation with eigenvalues $b_{1}^{3},b_{2},\ldots,b_{n}$. In turn, we conclude (38) by $x_{1}^{2}\geq\aleph^{2}$ for $x\not\in\Theta_{1,\aleph}$, and (37) by (35). ∎
Now, we are ready to prove a rather technical stability version of Theorem 1.3 that will lead to a stability version of Theorem 1.2.
Proposition 5.3.
Let $u\in S^{n-1}$ and $K\subseteq\mathbb{R}^{n}$ be a symmetric convex body, such that $K^{\circ}$ is in isotropic position with $|K^{\circ}|=\omega_{n}$, and let $\mathcal{P}$ be a Yao-Yao equipartition of the measure $d\mu(x)=\langle x,u\rangle^{2}\mathbf{1}_{K}(x)\,dx$ where $\mathcal{P}$ is centered at the origin, is based on $u^{\bot}$ and having axis $v\in S^{n-1}$. There exist explicit $C>0$, $R>n$ and $\varepsilon_{0}\in(0,1)$ depending only on $n$ satisfying that if
$\int_{K}\langle x,u\rangle^{2}\,dx\cdot\int_{K^{\circ}}\langle y,u\rangle^{2}\,dy\geq(1-\varepsilon)\left(\int_{B^{n}_{2}}\langle x,u\rangle^{2}\,dx\right)^{2}$ (39)
for $\varepsilon\in(0,\varepsilon_{0})$, then
$|\langle u,v\rangle|\geq{\color[rgb]{0,0,0}\sqrt{1-\varepsilon}}$ (40)
and for any $A\in\mathcal{P}$, one finds an $o$-symmetric ellipsoid $E_{A}$ (depending on $A$ and $K$) such that
$\big|\big((K\Delta E_{A})\cap A\big)\backslash\Theta_{u,\aleph}\big|\leq$ $\frac{C}{\aleph^{2}}\cdot\varepsilon^{\frac{1}{2}}\mbox{ \ for any $\aleph>0$,}$ (41)
$\frac{1}{R}\cdot B^{n}_{2}\subseteq$ $E_{A}\subseteq RB^{n}_{2}.$ (42)
Proof.
As $K^{\circ}$ is in isotropic position with $|K^{\circ}|=\omega_{n}$, we deduce from (34) that
$\frac{1}{n}\,B^{n}_{2}\subseteq K\subseteq nB^{n}_{2}.$ (43)
We may assume that $\langle u,v\rangle>0$. Let us consider again the linear transfrom $T|_{u^{\bot}}=\textnormal{Id}$, $T(v)=\langle u,v\rangle u$, which has the properties $\det T=1$ and
$T^{-1}(u)=\frac{v}{\langle u,v\rangle}.$ (44)
Recalling (22) and (24) one has
$\int_{\widetilde{K}}\langle y,u\rangle^{2}\,dy\cdot\int_{\widetilde{K}^{\circ}}\langle y,u\rangle^{2}\,dy=\frac{1}{\langle u,v\rangle{{}^{2}}}\int_{K}\langle x,u\rangle^{2}\,dx\cdot\int_{K^{\circ}}\langle x,v\rangle^{2}\,dx.$ (45)
As the left hand side of (45) is at most $\left(\int_{B^{n}_{2}}\langle x,u\rangle^{2}\,dx\right)^{2}$ by Theorem  1.3, we deduce the estimate $\langle u,v\rangle>{\color[rgb]{0,0,0}\sqrt{1-\varepsilon}}$ in (40) from (45) and the condition (39).
To prove (41), we observe that $\widetilde{\mathcal{P}}:=\{TA:\,A\in\mathcal{P}\}$ is a Yao-Yao equipartition of $\mathbb{R}^{n}$ with respect to the measure $d\tilde{\mu}(y)=\langle y,u\rangle^{2}\mathbf{1}_{\widetilde{K}}(y)\,dy$; moreover, $\widetilde{\mathcal{P}}$ is based on $u^{\bot}$ and has $u$ as its axis and the origin as its center. Let $A\in\mathcal{P}$. We may assume that $A=A^{\prime}+\mathbb{R}_{+}v$ where $A^{\prime}\in\mathcal{P}_{+}$ and $\mathcal{P}_{+}$ is the induced Yao-Yao partition of $u^{\bot}$, and $A^{\prime}={\rm pos}\{v_{1},\ldots,v_{n-1}\}$ for independent $v_{1},\ldots,v_{n-1}\in u^{\bot}\cap S^{n-1}$. In particular, $\widetilde{A}:=TA={\rm pos}\{A^{\prime},u\}\in\widetilde{\mathcal{P}}$.
It follows from the condition (39), (4) and (45) that
$(1-\varepsilon)2^{2n}\left(\int_{\mathbb{R}^{n}_{+}\cap B^{n}_{2}}\langle x,u\rangle^{2}\,dx\right)^{2}=$ $(1-\varepsilon)\left(\int_{B^{n}_{2}}\langle x,u\rangle^{2}\,dx\right)^{2}\leq\int_{\widetilde{K}^{\circ}}\langle x,u\rangle^{2}\,dx\cdot\int_{\widetilde{K}}\langle x,u\rangle^{2}\,dx$
$=$ $2^{n}\sum_{D\in\widetilde{\mathcal{P}}}\int_{D\cap\widetilde{K}}\langle x,u\rangle^{2}\,dx\cdot\int_{D^{\circ}\cap\widetilde{K}^{\circ}}\langle x,u\rangle^{2}\,dx.$
But, for any $D\in\widetilde{\mathcal{P}}$, as is obvious from the proof of (27) (simply observe that $D^{\circ}=-D^{*}$ and utilise the symmetry of $\widetilde{K}$),
$\int_{D\cap\widetilde{K}}\langle x,u\rangle^{2}\,dx\cdot\int_{D^{\circ}\cap\widetilde{K}^{\circ}}\langle x,u\rangle^{2}\,dx\leq\left(\int_{\mathbb{R}^{n}_{+}\cap B^{n}_{2}}\langle x,u\rangle^{2}\,dx\right)^{2},$
hence we deduce that
$\int_{\widetilde{A}\cap\widetilde{K}}\langle x,u\rangle^{2}\,dx\cdot\int_{\widetilde{A}^{\circ}\cap\widetilde{K}^{\circ}}\langle x,u\rangle^{2}\,dx\geq(1-2^{2n}\varepsilon)\left(\int_{\mathbb{R}^{n}_{+}\cap B^{n}_{2}}\langle x,u\rangle^{2}\,dx\right)^{2}.$ (46)
Next, denote $\Delta:=\det[v_{1},\dots,v_{n-1},v]$. Without loss of generality, we choose $u=e_{n}$. We consider the linear transform $S$ with $\det S=1$, where $Sv_{i}=\Delta^{\frac{1}{n-1}}e_{i}$ for $i=1,\ldots,n-1$ and $Su=u$. We also write $\Psi=T^{-1}S^{-1}$. We claim that
$\|\Psi\|_{\textnormal{op}}\leq M\qquad\textnormal{and}\qquad\|\Psi^{-1}\|_{\textnormal{op}}\leq M$ (47)
for some constant $M$ that depends only on $n$. For the first bound observe that if $x=\sum_{i=1}^{n}c_{i}e_{i}\in S^{n-1}$, then, for $\varepsilon\leq\frac{1}{2}$ (say), by (40) and (44),
$\|\Psi x\|_{2}=\left\|\sum_{i=1}^{n-1}c_{i}\Delta^{-\frac{1}{n-1}}v_{i}+c_{n}v/\langle u,v\rangle\right\|_{2}\leq\Delta^{-\frac{1}{n-1}}\sum_{i=1}^{n-1}|c_{i}|+\frac{|c_{n}|}{\langle u,v\rangle}\leq(n-1)\Delta^{-\frac{1}{n-1}}+2.$
But $\Delta$ is just the volume of the parallelepiped defined by the vectors $v_{1},\dots,v_{n-1},v$, which contains $A\cap B_{2}^{n}$. Thus $\|\Psi x\|_{2}\leq(n-1)V^{-\frac{1}{n-1}}+2$, where $V=|A\cap B_{2}^{n}|$. But $V$ has an explicit lower bound. Indeed, from (43) and from the facts that $\mathcal{P}$ is a Yao-Yao equipartition for $\mu$ and $\langle u,x\rangle^{2}\leq n^{2}$ for $x\in K$, we obtain
$V\geq\frac{|A\cap K|}{n^{n}}\geq\frac{1}{n^{2+n}}\int_{A\cap K}\langle u,x\rangle^{2}\,dx=\frac{1}{n^{2+n}2^{n}}\int_{K}\langle u,x\rangle^{2}\,dx\geq\frac{1}{n^{2+n}2^{n}}\int_{\frac{1}{2n}B^{n}_{2}}\langle u,x\rangle^{2}\,dx$ (48)
For the second bound assume that $x=\sum_{i=1}^{n-1}c_{i}v_{i}+c_{n}v\in S^{n-1}$ is a point such that $\|\Psi^{-1}x\|_{2}=\|\Delta^{\frac{1}{n-1}}\sum_{i=1}^{n-1}c_{i}e_{i}+c_{n}\langle u,v\rangle u\|_{2}$ may become arbitrarily large. Since $|\Delta|\leq 1$ and $|c_{n}|\leq 2$, without loss of generality $|c_{1}|$ becomes arbitrarily large, too. Take a $w\in S^{n-1}\cap\textnormal{span}(v_{2},\dots,v_{n-1},u)^{\perp}$. But then $|\langle c_{1}v_{1},w\rangle|\leq 1$ which implies that $\langle v_{1},w\rangle$ may come arbitrarily close to 0 hence $V$ may do the same, which contradicts (48), completing the proof of (47).
In turn, by the definition of the operator norm, (43) and (47) yield that the symmetric convex body $L:=S\widetilde{K}$ satisfies
$\frac{1}{nM}\cdot B^{n}_{2}\subseteq L\subseteq nM\cdot B^{n}_{2}.$ (49)
Recalling (28) and (30), (46) implies
$\int_{\mathbb{R}^{n}_{+}\cap L}\langle x,u\rangle^{2}\,dx\cdot\int_{\mathbb{R}^{n}_{+}\cap L^{\circ}}\langle x,u\rangle^{2}\,dx\geq$ $(1-2^{2n}\varepsilon)\left(\int_{\mathbb{R}^{n}_{+}\cap B^{n}_{2}}\langle x,u\rangle^{2}\,dx\right)^{2}.$ (50)
Given (50), we deduce from (37) and (38) in Proposition 5.2 the existence of a $\tilde{\theta}_{n}>1$ depending on $n$ and a positive definite diagonal matrix $\Phi$ (with respect to the orthonormal basis $e_{1},\ldots,e_{n}$) depending on $L$ and $\varepsilon$ such that
$\left|(\mathbb{R}^{n}_{+}\cap B^{n}_{2})\Delta\Phi(\mathbb{R}^{n}_{+}\cap L)\right|<$ $\tilde{\theta}_{n}\varepsilon^{\frac{n}{2(n+2)}};$ (51)
[…]
for any $\widetilde{\aleph}>0$ (remember that $e_{n}=u$).
Let $\Phi^{-1}e_{i}=t_{i}e_{i}$ for $i=1,\ldots,n$ and $t_{i}>0$. If $\varepsilon>0$ is small enough, then (49) shows that
$\color[rgb]{0,0,0}\frac{1}{2}<\det\Phi.$} (53)
It follows from (51), (52) and (53) that the unconditional ellipsoid $\widetilde{E}=\Phi^{-1}B^{n}_{2}$ satisfies
$\left|(\mathbb{R}^{n}_{+}\cap\widetilde{E})\Delta(\mathbb{R}^{n}_{+}\cap L)\right|<$ $2\tilde{\theta}_{n}\varepsilon^{\frac{n}{2(n+2)}};$ (54)
[…]
for any $\widetilde{\aleph}>0$. As $\frac{1}{{\color[rgb]{0,0,0}nM}}\cdot B^{n}_{2}\subseteq L$ by (49), we deduce from (54) that if $\varepsilon>0$ is small enough, then $t_{i}\geq\frac{1}{2{\color[rgb]{0,0,0}nM}}$ for $i=1,\ldots,n$, and hence (53) implies that $t_{j}\leq 2(2{\color[rgb]{0,0,0}nM})^{n-1}$ for $j=1,\ldots,n$ and
$\frac{1}{2{\color[rgb]{0,0,0}nM}}\cdot B^{n}_{2}\subseteq\widetilde{E}\subseteq 2(2{\color[rgb]{0,0,0}nM})^{n-1}\cdot B^{n}_{2}.$ (56)
It follows that $\Phi^{-1}\Theta_{u,\widetilde{\aleph}}\subseteq\Theta_{u,2(2{\color[rgb]{0,0,0}nM})^{n-1}\widetilde{\aleph}}$ for any $\widetilde{\aleph}>0$ in (55), or in other words, we have
$\left|(\mathbb{R}^{n}_{+}\cap\widetilde{E})\Delta(\mathbb{R}^{n}_{+}\cap L)\right|<$ $C\varepsilon^{\frac{n}{2(n+2)}};$ (57)
[…]
for arbitrary $\aleph>0$ where $C=2\tilde{\theta}_{n}\cdot 4(2{\color[rgb]{0,0,0}nM})^{2(n-1)}$ depends only on $n$. Now $\Psi$ satisfies that $\det\Psi=1$, $\Psi\Theta_{u,\aleph}=\Theta_{u,\aleph}$ and $A\cap K=\Psi(\mathbb{R}^{n}_{+}\cap L)$, thus we conclude (41) with $E_{A}=\Psi\widetilde{E}$.
Finally, (42) follows from (56) and (47), completing the proof of Proposition 5.3. ∎
6 Stability of Theorem 1.2
For a convex body $C\subseteq\mathbb{R}^{n}$, the inradius $r(C)$ denotes the maximal radius of any Euclidean balls contained in $C$.
Claim 6.1.
If $C\subseteq x+RB^{n}_{2}$ holds for a convex body $C\subseteq\mathbb{R}^{n}$, $x\in\mathbb{R}^{n}$ and $R>0$, then
$|C|\leq(n+1)\omega_{n-1}R^{n-1}r(C).$ (59)
Proof.
Let $z$ be the centroid of $C$, and let $\varrho>0$ be the maximal radius of any Euclidean balls contained in $K$ and centered at $z$, and hence there exists a $y\in(z+\varrho B^{n}_{2})\cap\partial C$. For the unit vector $u=\frac{y-z}{\|y-z\|}$, we have $h_{C-z}(-u)\leq nh_{C-z}(u)=n\varrho$ as $z$ is the centroid; therefore, $C$ is contained in the cylinder that is the intersection of $(x+RB^{n}_{2})+\mathbb{R}u$ and the strip between the hyperplanes $y+u^{\bot}$ and $y-(n+1)\varrho u+u^{\bot}$. ∎
During the proof of our stability statements concerning a symmetric convex body $K\subseteq\mathbb{R}^{n}$, we use various $o$-symmetric ellipsoids to approximate $K$ on a smaller part of $\mathbb{R}^{n}$, and the following lemma ensures that these ellipsoids are very close to each other.
Lemma 6.2.
For $n\geq 2$, $R>1$ and $s\in(0,\frac{1}{2})$, there exist explicit $\delta,C>0$ depending on $n,R,s$ such that if $\sigma\subseteq\mathbb{R}^{n}$ is a convex cone with $r(\sigma\cap B^{n}_{2})\geq s$, and the $o$-symmetric ellipsoids $E,E^{\prime}\subseteq\mathbb{R}^{n}$ satisy
$\frac{1}{R}\,B^{n}_{2}\subseteq E,E^{\prime}\subseteq RB^{n}_{2}\mbox{ \ and \ }|(E\Delta E^{\prime})\cap\sigma|\leq\delta,$ (60)
then
$|E\Delta E^{\prime}|\leq C\big|(E\Delta E^{\prime})\cap\sigma\big|.$
Proof.
During the proof of Lemma 6.2, we write $g\gg h$ or $h\ll g$ for some positive expressions $g$ and $h$ if $g\geq ch$ where $c>0$ is an explicit constant that depends only on $n,R,s$.
There exists a positive definite symmetric matrix $\Phi\in{\rm GL}(n)$ such that $\Phi E^{\prime}=B^{n}_{2}$, and the eigenvalues of $\Phi$ are between $\frac{1}{R}$ and $R$, and hence $R^{-n}\leq\det\Phi\leq R^{n}$. In particular, the $o$-symmetric ellipsoid $\widetilde{E}=\Phi E$ and the convex cone $\tilde{\sigma}=\Phi\sigma$ satisfy
$\frac{1}{R^{2}}\,B^{n}_{2}\subseteq$ $\widetilde{E}\subseteq R^{2}B^{n}_{2},$ (61)
$r(\tilde{\sigma}\cap B^{n}_{2})\geq$ $\frac{s}{R},$ (62)
and it is sufficient to find explicit $\delta,C>0$ depending on $n,R,s$ such that
$\mbox{if \ }\big|(\widetilde{E}\Delta B^{n}_{2})\cap\tilde{\sigma}\big|\leq\delta\cdot R^{n},\mbox{ \ then }|\widetilde{E}\Delta B^{n}_{2}|\leq\frac{C}{R^{n}}\cdot\big|(\widetilde{E}\Delta B^{n}_{2})\cap\sigma\big|.$ (63)
To prove (63), let $e_{1},\ldots,e_{n}$ be an orthonormal basis of $\mathbb{R}^{n}$ that form the principal directions of $\widetilde{E}$ in a way such that $a_{1}\leq\ldots\leq a_{n}$ hold for the half axes of $\widetilde{E}$ with $a_{i}e_{i}\in\partial\widetilde{E}$, $i=1,\ldots,n$ and $\langle e_{n},w\rangle\geq 0$. The core claim is that if (61) and (62) hold, and $\max_{i=1}^{n}\{|a_{i}-1|\}\geq\varepsilon$ for $\varepsilon\in(0,\frac{1}{2})$, then
$\left|(\widetilde{E}\Delta B^{n}_{2})\cap\tilde{\sigma}\right|\geq c\varepsilon$ (64)
where $c\in(0,1)$ depends on $n,R,s$.
For $u\in S^{n-1}$ and $\psi\in(0,\frac{\pi}{2}]$, we write $B(u,\psi)=\{x\in S^{n-1}:\langle x,u\rangle\geq\cos\psi\}$ to denote the spherical ball (cap) of center $u$ and geodesic radius $\psi$. It follows from the spherical triangle inequality that if $\eta\subseteq S^{n-1}$ is a great circle, and the spherical distance of $\eta$ from $u$ is at most $\frac{\psi}{2}$, then
[…]
According to (62), there exists $w\in S^{n-1}$ such that
$B(w,\alpha)\subseteq\tilde{\sigma}\cap S^{n-1}\mbox{ \ where }\sin\alpha=\frac{s}{R}.$ (66)
To verify the core claim (64), we distinguish two cases depending on $a_{n}$.
Case 1. $a_{n}\geq 1+\varepsilon$.
As $\langle e_{n},w\rangle\geq 0$, we have $w=v\cos\gamma+e_{n}\sin\gamma$ for a $v\in S^{n-1}\cap e_{n}^{\bot}$ and a $\gamma\in[0,\frac{\pi}{2}]$. We observe that any $z\in S^{n-1}\backslash\{\pm e_{n}\}$ can be written uniquely in the form $z=u\cos\varphi+e_{n}\sin\varphi$ for a $u\in S^{n-1}\cap e_{n}^{\bot}$ and a $\varphi\in(-\frac{\pi}{2},\frac{\pi}{2})$, and for $t\in(0,\frac{\pi}{4})$, we consider
$\Omega_{t}=\left\{u\cos\varphi+e_{n}\sin\varphi:\,u\in S^{n-1}\cap e_{n}^{\bot}\mbox{ and }\varphi\in\left(t,\frac{\pi}{2}-t\right)\right\}.$
For any bounded $\mathcal{H}^{n-1}$ measurable function $f:\Omega_{t}\to\mathbb{R}$, we have
$\int_{\Omega_{t}}f\,d\mathcal{H}^{n-1}=$ $\int_{S^{n-1}\cap e_{n}^{\bot}}\int_{t}^{\frac{\pi}{2}-t}f(u\cos\varphi+e_{n}\sin\varphi)(\cos\varphi)^{n-2}\,d\mathcal{H}^{1}(\varphi)d\mathcal{H}^{n-2}(u)$
$\geq$ $(\sin t)^{n-2}\int_{S^{n-1}\cap e_{n}^{\bot}}\int_{t}^{\frac{\pi}{2}-t}f(u\cos\varphi+e_{n}\sin\varphi)\,d\mathcal{H}^{1}(\varphi)d\mathcal{H}^{n-2}(u).$ (67)
We consider $w^{\prime}=e_{n}\cos(\gamma-\frac{3\alpha}{4})+v\sin(\gamma-\frac{3\alpha}{4})$ if $\gamma\geq\frac{\pi}{4}$ and $w^{\prime}=e_{n}\cos(\gamma+\frac{3\alpha}{4})+v\sin(\gamma+\frac{3\alpha}{4})$ if $\gamma<\frac{\pi}{4}$, and hence $B\left(w^{\prime},\frac{\alpha}{4}\right)\subseteq\widetilde{\sigma}\cap\Omega_{\alpha/2}$ by (66). We also consider
$\Xi=$ $B\left(w^{\prime},\frac{\alpha}{4}\right)\cap\left\{u\cos\varphi+e_{n}\sin\varphi:\,u\in B\left(v,\frac{\alpha}{8}\right)\cap e_{n}^{\bot}\mbox{ and }\varphi\in\left(0,\frac{\pi}{2}\right)\right\},$
$\xi_{u}=$ $\left\{\varphi\in\left(0,\frac{\pi}{2}\right):\,u\cos\varphi+e_{n}\sin\varphi\in\Xi\right\}\mbox{ for a $u\in B\left(v,\frac{\alpha}{8}\right)\cap e_{n}^{\bot}$.}$
In particular, $\Xi$ and $\eta_{u}$ satisfy via (65) that
$\Xi\subseteq$ $\widetilde{\sigma}\cap\Omega_{\alpha/2},$ (68)
$\xi_{u}$ is an interval of length at least $\frac{\alpha}{4}$ for any fixed $u\in B\left(v,\frac{\alpha}{8}\right)\cap e_{n}^{\bot}$. (69)
Writing $\varrho_{\widetilde{E}}$ to denote the radial function of $\widetilde{E}$, we deduce from (66), $\sin\frac{\alpha}{2}\geq\frac{s}{2R}$, (67) and (68) that
[…]
$\gg$ $\int_{B\left(v,\frac{\alpha}{8}\right)\cap e_{n}^{\bot}}\int_{\xi_{u}}\left|1-\varrho_{\widetilde{E}}(u\cos\varphi+e_{n}\sin\varphi)^{n}\right|\,d\mathcal{H}^{1}(\varphi)d\mathcal{H}^{n-2}(u).$
In particular, our core claim (64) follows in Case 1 if for any $u\in B\left(v,\frac{\alpha_{n}}{8}\right)$, we have
$\int_{\xi_{u}}\left|1-\varrho_{\widetilde{E}}(u\cos\varphi+e_{n}\sin\varphi)^{n}\right|\,d\varphi\gg\varepsilon.$ (70)
To prove (70), let $u\in B\left(v,\frac{\alpha}{8}\right)\cap e_{n}^{\bot}$. Now, $\widetilde{E}\cap{\rm lin}\{u,e_{n}\}$ is an ellipse with principle axes $e_{n}$ and $u$, and corresponding half axes $a_{n}$ and $\varrho_{\widetilde{E}}(u)\leq a_{n}$, and hence
$\varrho_{\widetilde{E}}(u\cos\varphi+e_{n}\sin\varphi)^{2}=\frac{a^{2}_{n}\varrho_{\widetilde{E}}(u)^{2}}{a_{n}^{2}(\cos\varphi)^{2}+\varrho_{\widetilde{E}}(u)^{2}(\sin\varphi)^{2}}=\frac{a^{2}_{n}\varrho_{\widetilde{E}}(u)^{2}}{(a_{n}^{2}-\varrho_{\widetilde{E}}(u)^{2})(\cos\varphi)^{2}+\varrho_{\widetilde{E}}(u)^{2}}.$ (71)
It follows that $\varrho_{\widetilde{E}}(u\cos\varphi+e_{n}\sin\varphi)$ is an increasing function of $\varphi$. The segment $\xi_{u}$ is divided by its midpoint $z_{u}$ into two segments $\xi_{-}$ and $\xi_{+}$ of length at least $\frac{\alpha}{8}$ according to (69) where $\varphi\leq z_{u}$ if $\varphi\in\xi_{-}$ and $\varphi\geq z_{u}$ if $\varphi\in\xi_{+}$. We consider three cases:
Case 1.1.
If $\varrho_{\widetilde{E}}(z_{u})\geq 1+\frac{\varepsilon}{2}$, then $\varrho_{\widetilde{E}}(u\cos\varphi+e_{n}\sin\varphi)^{n}\geq(1+\frac{\varepsilon}{2})^{n}$ for $\varphi\in\xi_{+}$, and the integral over $\xi_{+}$ yields (70).
Case 1.2.
If $1-\varepsilon<\varrho_{\widetilde{E}}(z_{u})<1+\frac{\varepsilon}{2}$, then (71) yields that for $\varphi\in\xi_{+}$, we have $\sin 2\varphi\geq\sin\alpha_{n}$ and
$\frac{\partial}{\partial\varphi}\varrho_{\widetilde{E}}(u\cos\varphi+e_{n}\sin\varphi)^{n}=\frac{n\varrho_{\widetilde{E}}(u\cos\varphi+e_{n}\sin\varphi)^{n}}{2}\cdot\frac{(a_{n}^{2}-\varrho_{\widetilde{E}}(u)^{2})\sin 2\varphi}{a_{n}^{2}(\cos\varphi)^{2}+\varrho_{\widetilde{E}}(u)^{2}(\sin\varphi)^{2}}\gg\varepsilon$
(which estimate works even if $a_{n}$ is large), thus again the integral over $\xi_{+}$ implies (70).
Case 1.3.
If $\varrho_{\widetilde{E}}(z_{u})\leq 1-\varepsilon$, then $\varrho_{\widetilde{E}}(u\cos\varphi+e_{n}\sin\varphi)^{n}\leq(1-\frac{\varepsilon}{2})^{n}$ for $\varphi\in\xi_{-}$, and the integral over $\xi_{-}$ yields (70).
Since we have verified (70) in all sub-cases, we conclude the core claim (64) in Case 1.
Case 2. $a_{n}<1+\varepsilon$, and hence $a_{1}\leq 1-\varepsilon$.
In this case, the argument verifying (64) is very similar to the one in Case 1 only replacing $a_{n}$ by $a_{1}$; therefore, we do not present it in detail. In this case, $w=v\cos\gamma+e_{1}\sin\gamma$ for a $v\in S^{n-1}\cap e_{1}^{\bot}$ and a $\gamma\in[0,\frac{\pi}{2}]$, we have
$\Omega_{t}=\left\{u\cos\varphi+e_{1}\sin\varphi:\,u\in S^{n-1}\cap e_{1}^{\bot}\mbox{ and }\varphi\in\left(t,\frac{\pi}{2}-t\right)\right\}$
for $t\in(0,\frac{\pi}{4})$, and
$\Xi=\left\{u\cos\varphi+e_{1}\sin\varphi\in B\left(w^{\prime},\frac{\alpha}{4}\right):\,u\in B\left(v,\frac{\alpha}{8}\right)\cap e_{1}^{\bot}\mbox{ and }\varphi\in\left(0,\frac{\pi}{2}\right)\right\}.$
For $u\in B\left(v,\frac{\alpha_{n}}{8}\right)\cap e_{1}^{\bot}$, we have $a_{1}\leq\varrho_{\widetilde{E}}(z_{u})\leq a_{n}<1+\varepsilon$, and hence only the analogues of the sub-cases Case 1.1 and Case 1.2 occur when verifying the analogue of (70). Therefore, we have completed the proof of the core claim (64).
Finally, having the core claim (64) at hand, we set $\delta=\frac{c}{2n}$ for the $c$ in (64). If $\left|(E\Delta B^{n}_{2})\cap\tilde{\sigma}\right|\leq\delta$, then $\varepsilon=\max_{i=1}^{n}\{|a_{i}-1|\}\leq\frac{1}{2n}$, and hence
[…]
as $e^{t}\leq 1+2t$ for $t\in(0,1)$, verifying (63) with $C=\frac{3nR^{n}\omega_{n}}{c}$, and in turn proving Lemma 6.2. ∎
We are ready to verify the stability version Theorem 6.3, - equivalent to Theorem 1.5 - of Theorem 1.2.
Theorem 6.3.
There exist $C_{n}>1$ and $\varepsilon_{0}\in(0,1)$ depending on $n\geq 2$ such that if $\varepsilon\in(0,1)$ and symmetric convex body $K\subseteq\mathbb{R}^{n}$ satisfy
$\int_{K}\int_{K^{\circ}}\langle x,y\rangle^{2}\,dydx>\left(1-\varepsilon\right)\int_{B^{n}_{2}}\int_{B^{n}_{2}}\langle x,y\rangle^{2}\,dydx,$ (72)
then
$|E\Delta K|\leq C_{n}\sqrt{\varepsilon}$ (73)
for some $o$-symmetric ellipsoid $E\subseteq\mathbb{R}^{n}$.
Proof.
We may assume that $K^{\circ}$ is in isotropic position and $|K^{\circ}|=\omega_{n}$, and hence (34) yields that
$\frac{1}{n}\,B^{n}_{2}\subseteq K\subseteq nB^{n}_{2}.$ (74)
Let us discuss some consequences of the conditions (72) and that $K^{\circ}$ is in isotropic position before we start the actual proof of Theorem 6.3. For any $u\in S^{n-1}$, we claim that
$\int_{K}\langle x,u\rangle^{2}\,dx\cdot\int_{K^{\circ}}\langle y,u\rangle^{2}\,dy\geq(1-n\varepsilon)\left(\int_{B^{n}_{2}}\langle x,u\rangle^{2}\,dx\right)^{2}.$ (75)
To prove (75), we may assume $u=e_{n}$. We deduce from and (72) that
$\sum_{i=1}^{n}\int_{K}\langle x,e_{i}\rangle^{2}\,dx\cdot\int_{K^{\circ}}\langle y,e_{i}\rangle^{2}\,dy\geq(1-\varepsilon)\int_{B^{n}_{2}}\int_{B^{n}_{2}}\langle x,y\rangle^{2}\,dydx=(1-\varepsilon)n\left(\int_{B^{n}_{2}}\langle x,u\rangle^{2}\,dx\right)^{2}.$
As $\int_{K}\langle x,e_{i}\rangle^{2}\,dx\cdot\int_{K^{\circ}}\langle y,e_{i}\rangle^{2}\,dy\leq\left(\int_{B^{n}_{2}}\langle x,u\rangle^{2}\,dx\right)^{2}$ holds for $i=1,\ldots,n$ by Theorem 1.3, we conclude (75).
Next, we discuss some properties of Yao-Yao partitions. For any $u\in S^{n-1}$, we consider the even measure $d\mu(x)=\langle x,u\rangle^{2}\mathbf{1}_{K}(x)\,dx$, and a Yao-Yao equipartition $\mathcal{P}$ of $\mathbb{R}^{n}$ for $\mu$ based on $u^{\bot}$, centered at the origin and having an axis $v\in S^{n-1}$ satisfying that $\langle u,v\rangle>0$. It follows from (75) and Proposition 5.3 that there exist $\widetilde{C}>0$, $R>n$ and $\tilde{\varepsilon}_{0}\in(0,1)$ depending only on $n$ such that if $\varepsilon<\tilde{\varepsilon}_{0}$, then $\langle u,v\rangle\geq 1-n\varepsilon$, and hence the Law of Cosine yields
$\|v-u\|_{2}\leq\sqrt{2n\varepsilon};$ (76)
moreover, for any $A\in\mathcal{P}$, one finds an $o$-symmetric ellipsoid $E_{A}$ (depending on $A$ and $K$) with the properties
$\big|\big((K\Delta E_{A})\cap A\big)\backslash\Theta_{u,\aleph}\big|\leq$ $\frac{\widetilde{C}}{\aleph^{2}}\cdot\varepsilon^{\frac{1}{2}}\mbox{ \ for any $\aleph>0$,}$ (77)
$\frac{1}{R}\cdot B^{n}_{2}\subseteq$ $E_{A}\subseteq RB^{n}_{2}.$ (78)
In addition, we claim that there exists an $\tilde{r}>0$ depending only on $n$ such that if $A\in\mathcal{P}$, then
$r(A\cap K)\geq\tilde{r}\mbox{ \ for any }A\in\mathcal{P}.$ (79)
To prove (79), note that $\langle u,x\rangle^{2}\leq n^{2}$ holds for any $x\in K$ by (74), and hence, exactly as in the proof of (48), $|A\cap K|\geq\delta_{n}$ for some computable constant $\delta_{n}$ that depends only on $n$. Therefore, Claim 6.1 yields (79).
For a fixed $u_{0}\in S^{n-1}$, we consider the measure $d\mu_{0}(x)=\langle x,u_{0}\rangle^{2}\mathbf{1}_{K}(x)\,dx$, and a Yao-Yao equipartition $\mathcal{P}_{0}$ of $\mathbb{R}^{n}$ for $\mu_{0}$ based on $u_{0}^{\bot}$ and centered at the origin. We also fix an $A_{0}\in\mathcal{P}_{0}$ with $A_{0}\subseteq\{x\in\mathbb{R}^{n}:\langle u_{0},x\rangle\geq 0\}$. It follows from (79) that there exists a $\tilde{z}_{0}\in A_{0}\cap K$ such that $\tilde{z}_{0}+\tilde{r}B^{n}_{2}\subseteq A_{0}\cap K$, and hence one finds a $z_{0}\in\tilde{z}_{0}+\tilde{r}B^{n}_{2}$ such that (cf. (74))
$z_{0}+\frac{\tilde{r}}{2}\,B^{n}_{2}\subseteq A_{0}\cap K\subseteq nB^{n}_{2}\mbox{ \ and \ }\langle u_{0},y\rangle\geq\tilde{r}\mbox{ \ for }y\in z_{0}+\frac{\tilde{r}}{2}\,B^{n}_{2}.$ (80)
We apply (77) and (78) with $\aleph=\frac{\tilde{r}}{nR}$ to $A_{0}\in\mathcal{P}_{0}$, and deduce that there exists an $o$-symmetric ellipsoid $E_{0}$ such that (cf. (74) and $R\geq n$)
$\big|\big((K\Delta E_{0})\cap A_{0}\big)\backslash{\color[rgb]{0,0,0}\Theta_{u_{0},\frac{\tilde{r}}{nR}}}\big|\leq$ $C_{0}\cdot\varepsilon^{\frac{1}{2}}$ (81)
$\frac{1}{R}\cdot B^{n}_{2}\subseteq$ $E_{0}\cap K\subseteq E_{0}\subseteq RB^{n}_{2}$ (82)
where $C_{0}>0$ depends only on $n$.
We consider the convex cone
$\sigma_{0}={\rm pos}\left(z_{0}+\frac{\tilde{r}}{2}\,B^{n}_{2}\right)\subseteq A_{0}.$ (83)
If $x\in\sigma_{0}\cap\Theta_{u,\frac{\tilde{r}}{nR}}$, then $x=ty$ for a $y\in z_{0}+\frac{\tilde{r}}{2}\,B^{n}_{2}$ and $t\geq 0$, and hence $t\leq\frac{1}{nR}$ by (80). As $\|y\|_{2}\leq n$ also by (80), we deduce that $x\in\frac{1}{R}\cdot B^{n}_{2}\subseteq E_{0}\cap K$ by (82), thus (81) yields that
$\big|(K\Delta E_{0})\cap\sigma_{0}\big|\leq C_{0}\cdot\varepsilon^{\frac{1}{2}}.$ (84)
Let $w_{1},\ldots,w_{n}\in S^{n-1}\cap z_{0}^{\bot}$ be the vertices of a regular (n1)(n-1)-simplex, and let $y_{i}=z_{0}+\frac{\tilde{r}}{4}\,w_{i}$ and $u_{i}=y_{i}/\|y_{i}\|_{2}$ for $i=1,\ldots,n$. Here, $y_{i}=t_{0}u_{i}$ for $t_{0}=\|y_{i}\|_{2}$, $i=1,\ldots,n$, and
$n\geq t_{0}>\|z_{0}\|_{2}\geq\frac{3}{2}\,\tilde{r}$
hold by (80). We claim that there exists $\tau>0$ depending only on $n$ such that
$t_{0}u_{i}+\frac{\tilde{r}}{4}\,B^{n}_{2}\subseteq$ $\sigma_{0}\cap K$ $\mbox{for }i=1,\ldots,n,$ (85)
$\langle u_{i},y\rangle\geq$ $\tilde{r}$ $\mbox{for }y\in t_{0}u_{i}+\frac{\tilde{r}}{4}\,B^{n}_{2},$ (86)
$\bigcap_{i=1}^{n}\Theta_{u_{i},\aleph}\subseteq$ $\tau\aleph\,B^{n}_{2}$ $\mbox{for any }\aleph>0.$ (87)
Here (85) directly follows from (80) and (83), and (86) follows as $t_{0}=\|y_{i}\|_{2}\geq\frac{3}{2}\,\tilde{r}$.
For (87), we observe that the volume of the simplex $\Sigma={\rm conv}\{o,y_{1},\ldots,y_{n}\}$ is at least $\tau_{0}$ for a $\tau_{0}>0$ depending only on $n$ by construction and $\|z_{0}\|_{2}\geq\frac{3}{2}\,\tilde{r}$, and $\Sigma\subseteq nB^{n}_{2}$ by (85). Thus, $r(\Sigma)\geq\tilde{\tau}$ holds for a $\tilde{\tau}>0$ depending only on $n$ by Claim 6.1, and in turn $r(\Sigma_{0})\geq\tilde{\tau}/n$ for the simplex $\Sigma_{0}={\rm conv}\{o,u_{1},\ldots,u_{n}\}$ by $t_{0}\leq n$. Now, if $p_{1},\ldots,p_{n}$ form the dual basis for $u_{1},\ldots,u_{n}$; namely, $\langle p_{i},u_{i}\rangle=1$ and $\langle p_{i},u_{j}\rangle=0$ for $j\neq i$, then $r(\Sigma_{0})\geq\tilde{\tau}/n$ implies that $\|p_{i}\|_{2}\leq n/\tilde{\tau}$ for $i=1,\ldots,n$, and hence if $x\in\bigcap_{i=1}^{n}\Theta_{u_{i},\aleph}$, then
$\|x\|_{2}=\left\|\sum_{i=1}^{n}\langle x,u_{i}\rangle p_{i}\right\|_{2}\leq\frac{n^{2}}{\tilde{\tau}}\cdot\aleph$,
proving (87) with $\tau=n^{2}/\tilde{\tau}$.
Our core claim is that there exists $\widetilde{\aleph}>0$ depending only on $n$ such that $\widetilde{\aleph}\leq\frac{1}{\tau R}$, and for any $i=1,\ldots,n$, we have
$\big|(K\Delta E_{0})\backslash\Theta_{u_{i},\widetilde{\aleph}}\big|\leq\widetilde{C}_{0}\cdot\varepsilon^{\frac{1}{2}}$ (88)
where $\widetilde{C}_{0}>0$ depends only on $n$.
To prove (88), we consider a Yao-Yao equipartition $\mathcal{P}_{i}$ of $\mathbb{R}^{n}$ for the measure $d\mu_{i}(x)=\langle x,u_{i}\rangle^{2}\mathbf{1}_{K}(x)\,dx$ based on $u_{i}^{\bot}$, centered at the origin and having an axis $v_{i}\in S^{n-1}$ satisfying that $\langle u_{i},v_{i}\rangle>0$. We deduce from $t_{0}\leq n$ and (76) that $\|t_{0}v_{i}-t_{0}u_{i}\|_{2}\leq n\sqrt{2n\varepsilon}$. Therefore, if $\varepsilon>0$ is small enough, then (85) and (86) yield that
$t_{0}v_{i}+\frac{\tilde{r}}{8}\,B^{n}_{2}\subseteq K\cap\sigma_{0}\mbox{ \ and \ }\langle u_{i},y\rangle\geq\tilde{r}\mbox{ \ for }y\in t_{0}v_{i}+\frac{\tilde{r}}{8}\,B^{n}_{2}.$ (89)
Let $A_{1},\ldots,A_{2^{n-1}}$ be the elements of $\mathcal{P}_{i}$ containing $v_{i}$, and hence $t_{0}v_{i}\in A_{j}\cap K$ for each $A_{j}$. For any $j=1,\ldots,2^{n-1}$, there exists a $\tilde{z}_{j}\in A_{j}\cap K$ such that $\tilde{z}_{j}+\tilde{r}B^{n}_{2}\subseteq A_{j}\cap K$ by (79). As $\tilde{z}_{j}+\tilde{r}B^{n}_{2}\subseteq t_{0}v_{i}+2nB^{n}_{2}$ according to (74), (89) implies that $z_{j}=t_{0}v_{i}+\frac{\tilde{r}}{16n}(\tilde{z}_{j}-t_{0}v_{j})$ satisfies
$z_{j}+\frac{\tilde{r}^{2}}{16n}\,B^{n}_{2}\subseteq K\cap\sigma_{0}\cap A_{j}\mbox{ \ and \ }\langle u_{i},y\rangle\geq\tilde{r}\mbox{ \ for }y\in z_{j}+\frac{\tilde{r}^{2}}{16n}\,B^{n}_{2}.$ (90)
We apply (77) and (78) to $A_{j}\in\mathcal{P}_{i}$ with
$\aleph=\widetilde{\aleph}:=\min\left\{\frac{\tilde{r}}{nR},\frac{1}{\tau R}\right\},$
and deduce that there exists an $o$-symmetric ellipsoid $E_{j}$ such that (cf. (74) and $R\geq n$)
[…]
$\frac{1}{R}\cdot B^{n}_{2}\subseteq$ $E_{j}\cap K\subseteq E_{j}\subseteq RB^{n}_{2}$ (92)
where $C^{\prime}>0$ depends only on $n$.
We consider the convex cone
$\sigma_{j}={\rm pos}\left(z_{j}+\frac{\tilde{r}}{8}\,B^{n}_{2}\right)\subseteq A_{j}\cap\sigma_{0}.$ (93)
If $x\in\sigma_{j}\cap\Theta_{u_{i},\widetilde{\aleph}}$, then $x=ty$ for a $y\in z_{j}+\frac{\tilde{r}}{8}\,B^{n}_{2}$ and $t\geq 0$, and hence $t\leq\frac{1}{nR}$ by (90). As $\|y\|_{2}\leq n$ also by (90), we deduce that $x\in\frac{1}{R}\cdot B^{n}_{2}\subseteq E_{j}\cap K$ by (92), thus (91) yields that
$\big|(K\Delta E_{j})\cap\sigma_{j}\big|\leq C^{\prime}\cdot\varepsilon^{\frac{1}{2}}.$ (94)
Comparing (94) to (84), and $\sigma_{j}\subseteq\sigma_{0}$ (cf. (93)) yield that
$\big|(E_{0}\Delta E_{j})\cap\sigma_{j}\big|\leq(C^{\prime}+C_{0})\cdot\varepsilon^{\frac{1}{2}}.$ (95)
In turn, we deduce from Lemma 6.2, (82), (92) and (93) that $|E_{0}\Delta E_{j}|\leq C^{*}\cdot\varepsilon^{\frac{1}{2}}$ for a $C^{*}>0$ depending only on $n$, and hence (91) yields
$\big|\big((K\Delta E_{0})\cap A_{j}\big)\backslash\Theta_{u_{i},\widetilde{\aleph}}\big|\leq(C^{\prime}+C^{*})\cdot\varepsilon^{\frac{1}{2}}.$ (96)
As $\bigcup_{j=1}^{2^{n-1}}A_{j}=\{x\in\mathbb{R}^{n}:\langle x,u_{i}\rangle\geq 0\}$, and $K$ and $E_{0}$ are $o$-symmetric, it follows from (96) that
$\big|(K\Delta E_{0})\backslash\Theta_{u_{i},\widetilde{\aleph}}\big|\leq 2^{n}(C^{\prime}+C^{*})\cdot\varepsilon^{\frac{1}{2}},$ (97)
proving the core claim (88) with $\widetilde{C}_{0}=2^{n}(C^{\prime}+C^{*})$ for any $i=1,\ldots,n$.
Now, $\bigcap_{i=1}^{n}\Theta_{u_{i},\widetilde{\aleph}}\subseteq E_{0}\cap K$ by $\widetilde{\aleph}\leq\frac{1}{\tau R}$, (87) and $\frac{1}{R}\,B^{n}_{2}\subseteq E_{0}\cap K$; therefore, our core claim (88) implies $|K\Delta E_{0}|\leq n\widetilde{C}_{0}\varepsilon^{\frac{1}{2}}$, completing the proof of Theorem 6.3. ∎
7 Proof of Theorem 1.6 and Collorary 1.7
Comparing a symmetric convex body $L\subseteq\mathbb{R}^{n}$ with the centered ball of the same volume shows that
$|L|^{\frac{n+2}{n}}\leq\gamma_{n}\int_{L}\|x\|^{2}\,dx\mbox{ \ \ for \ }\gamma_{n}=\frac{(n+2)\omega_{n}^{\frac{2}{n}}}{n},$ (98)
where equality holds if and only if $L$ is a Euclidean ball. We also recall that if $L$ is in isotropic position, then
$\int_{L}\|x\|^{2}\,dx=n\int_{L}\langle x,u\rangle^{2}\,dx\mbox{ \ \ for any \ }u\in S^{n-1}.$ (99)
If $K$ is a symmetric convex body such that $K^{\circ}$ is in isotropic position, then it follows from first (98), then applying $\|x\|^{2}=\sum_{i=1}^{n}\langle x,e_{i}\rangle^{2}$ to $K$ and (99) to $L=K^{\circ}$ and $u=e_{1},\ldots,e_{n}$, and after that (33), and finally Theorem 1.2, and then reversing the path, we deduce that
$\left(|K|\cdot|K^{\circ}|\right)^{\frac{n+2}{n}}\leq$ $\gamma^{2}_{n}\int_{K}\|x\|^{2}\,dx\cdot\int_{K^{\circ}}\|x\|^{2}\,dx=n\gamma^{2}_{n}\sum_{i=1}^{n}\int_{K}\langle x,e_{i}\rangle^{2}\,dx\cdot\int_{K^{\circ}}\langle x,e_{i}\rangle^{2}\,dx$ (100)
$=$ $n\gamma^{2}_{n}\int_{K}\int_{K^{\circ}}\langle x,y\rangle^{2}\,dx\,dy\leq n\gamma^{2}_{n}\int_{B_{2}^{n}}\int_{B^{n}_{2}}\langle x,y\rangle^{2}\,dx\,dy$
$=$ $\gamma^{2}_{n}\int_{B_{2}^{n}}\|x\|^{2}\,dx\cdot\int_{B_{2}^{n}}\|x\|^{2}\,dx=\left(|B_{2}^{n}|\cdot|B_{2}^{n}|\right)^{\frac{n+2}{n}}.$ (101)
Comparing (100) and (101) shows how Theorem 1.2 yields the Blaschke-Santaló inequality (2), but we also deduce the inequality (8) for the moment product. Actually, combining (100) and (101) with Theorem 1.5 directly yields the stability version Theorem 1.6 of the Blaschke-Santaló inequality.
For Corollary 1.7 concerning the affine surface area, we note that (cf. Lutwak [26])
$\Omega(K)\leq$ $n|K^{*}|^{\frac{1}{n+1}}|K|^{\frac{n}{n+1}}$
In particular, Theorem 1.6 yields Corollary 1.7.
Finally, we show that the error term in Theorem 1.5, Theorem 1.6 and Corollary 1.7 is of optimal order. For symmetric convex bodies $K,M\subseteq\mathbb{R}^{n}$, the stability estimates in the statements above can be expressed in terms of the linear invariant quantity
$\widetilde{A}(K,M)=\min_{\Phi\in{\rm GL}(n)}A(K,\Phi M)$
where it is easy to see that we can write minimum in the definition. For example, the stability version Theorem 1.6 of the Blaschke-Santaló inequality can be written in the form
$|K|\cdot|K^{\circ}|\leq\left(1-\theta_{n}\cdot\widetilde{A}(K,B^{n}_{2})^{2}\right)\omega_{n}^{2}$ (102)
where $\theta_{n}>0$ depends only on $n$.
Example 7.1 (Optimality of Theorem 1.5, Theorem 1.6 and Corollary 1.7).
According to the argument above in Section 7, it is sufficient to verify the optimality of the exponent $2$ of $\widetilde{A}(K,B^{n}_{2})$ in the error term for the stability version (102) of the Blaschke-Santaló inequality. In particular, we prove that if $t\in(-\eta,\eta)$ for some $\eta>0$ depending on $n$, then there exists a family of convex bodies $K_{t}\subseteq\mathbb{R}^{n}$ and constants $c_{0},C_{0}>0$ depending only on $n$ such that
(i) $|K_{t}|=\omega_{n}$ and $|K_{t}\Delta B^{n}_{2}|\leq C_{0}|t|$;
(ii) $K_{0}=B^{n}_{2}$;
(iii) $|K_{t}\Delta E|\geq c_{0}|t|$ for any $o$-symmetric ellipsoid $E\subseteq\mathbb{R}^{n}$ with $|E|=\omega_{n}$;
(iv) $|K_{t}^{\circ}|\geq(1-C_{0}t^{2})\omega_{n}$.
According to the properties (i)-(iv), the family $K_{t}$ satisfies that $\widetilde{A}(K_{t},B^{n}_{2})$ tends to zero as $t$ tends to zero, and
$|K_{t}|\cdot|K_{t}^{\circ}|\geq\left(1-C\cdot\widetilde{A}(K_{t},B^{n}_{2})^{2}\right)\omega_{n}^{2}\mbox{ \ for \ }C=\frac{C_{0}}{c_{0}\omega_{n}^{1/n}},$
verifying the optimality of the error term in the stability version (102) of the Blaschke-Santaló inequality.
To construct $K_{t}$ if $|t|$ is small, let $e_{1},\ldots,e_{n}$ be an orthonormal basis of $\mathbb{R}^{n}$, and let $u_{0}=\frac{e_{1}+\ldots+e_{n}}{\sqrt{n}}\in S^{n-1}$, and hence
$\|u_{0}-e_{i}\|>\frac{1}{2}\mbox{ \ for }i=1,\ldots,n.$
We fix a $C^{\infty}$ function $\tilde{\varphi}:S^{n-1}\to[0,1]$ such that for $u\in S^{n-1}$, we have
$\tilde{\varphi}(u)=$ $0\mbox{ \ if }\|u-u_{0}\|_{2}\geq\frac{1}{4}$
$\tilde{\varphi}(u)=$ $1\mbox{ \ if }\|u-u_{0}\|_{2}\leq\frac{1}{8},$
and consider the even $C^{\infty}$ function $\varphi:S^{n-1}\to[0,1]$ defined by
$\varphi(u)=\tilde{\varphi}(u)+\tilde{\varphi}(-u)\mbox{ \ for }u\in S^{n-1}.$
Now, there exists $\eta\in(0,1)$ such that if $t\in[-\eta,\eta]$, then the even $C^{\infty}$ function $h_{t}=1+t\varphi$ on $S^{n-1}$ satisfies that the symmetric matrix $\nabla^{2}h_{t}+h_{t}I$ is positive definite where $\nabla^{2}h_{t}$ is the spherical Hessian and $\nabla h$ is the spherical gradient with respect to a moving frame, and hence there exists a symmetric convex body $\widetilde{K}_{t}\subseteq\mathbb{R}^{n}$ with $C^{\infty}_{+}$ boundary such that for any $u\in S^{n-1}$, we have
[…]
We also consider the convex cones $\sigma={\rm pos}\{u\in S^{n-1}:\|u-e_{1}\|_{2}\leq\frac{1}{4}\}$ and $\sigma_{0}={\rm pos}\{u\in S^{n-1}:\|u-u_{0}\|_{2}\leq\frac{1}{4}\}$. Since
$|\widetilde{K}_{t}|=$ $\frac{1}{n}\int_{S^{n-1}}h_{t}\det(\nabla^{2}h_{t}+h_{t}I)\,d\mathcal{H}^{n-1},$
$|\widetilde{K}_{t}^{*}|=$ $\frac{1}{n}\int_{S^{n-1}}h_{t}^{-n}\,d\mathcal{H}^{n-1},$
we deduce that $|\widetilde{K}_{t}|$ and $|\widetilde{K}_{t}^{*}|$ are $C^{\infty}$ functions of $t\in(-\eta,\eta)$; moreover,
$(1-|t|)B^{n}_{2}\subseteq$ $\widetilde{K}_{t}\subseteq(1+|t|)B^{n}_{2},$
$h_{\widetilde{K}_{t}}(u)=\varrho_{\widetilde{K}_{t}}(u)=$ $1\mbox{ \ if }u\in\sigma\cap S^{n-1},$
$h_{\widetilde{K}_{t}}(u)=\varrho_{\widetilde{K}_{t}}(u)=$ $1+t\mbox{ \ if }u\in\sigma_{0}\cap S^{n-1}.$
After possibly choosing $\eta>0$ smaller, the renormalized body
$K_{t}=\lambda_{t}\cdot\widetilde{K}_{t}\mbox{ \ for }\lambda_{t}=\frac{\omega_{n}^{\frac{1}{n}}}{|\widetilde{K}_{t}|^{\frac{1}{n}}}\in\left(\frac{3}{4},\frac{4}{3}\right)$
satisfies that $|K_{t}|=\omega_{n}$, $\frac{1}{2}B^{n}_{2}\subseteq K_{t}\subseteq 2B^{n}_{2}$; moreover, $\lambda_{t}$ and $|K_{t}^{*}|$ are $C^{\infty}$ functions of $t\in(-\eta,\eta)$, and hence $\lambda_{t}=1+O(|t|)$, and
$\left(1-O(|t|)\right)B^{n}_{2}\subseteq$ $K_{t}\subseteq\left(1+O(|t|)\right)B^{n}_{2},$
$h_{K_{t}}(u)=\varrho_{K_{t}}(u)=$ $\lambda_{t}\mbox{ \ if }u\in\sigma\cap S^{n-1},$ (103)
$h_{K_{t}}(u)=\varrho_{K_{t}}(u)=$ $(1+t)\lambda_{t}\mbox{ \ if }u\in\sigma_{0}\cap S^{n-1}.$ (104)
We are ready to prove that $K_{t}$ satisfies (i)-(iv). Here (i) and (ii) readily holds. For (iv), we observe that that $f(t)=|K_{t}^{*}|$ is a $C^{\infty}$ function of $t$ near 0, and the Blaschke-Santaló inequality yields that $f(t)$ has a maximum at $t=0$, and hence $f^{\prime}(t)=0$. In turn, the Taylor formula $f(t)=f(0)+f^{\prime}(0)t+\frac{1}{2}\,f"(\xi)t^{2}$ where $|\xi|\leq|t|$ yields that $f(t)\geq f(0)-C_{0}t^{2}$ for a constant $C_{0}>0$ independent of $t$, verifying (iv).
Finally, we prove (iii) indirectly; therefore, we suppose that there exists a sequence $t_{k}\neq 0$ tending to zero as $k$ tends to infinity and a sequence of $o$-symmetric ellipsoids $E_{t_{k}}$ such that $|E_{t_{k}}|=\omega_{n}$ and
$\left|K_{t_{k}}\Delta E_{t_{k}}\right|=o(|t_{k}|)$
as $k$ tends to infinity, and seek a contradiction. We deduce from (103) and (104) that
$\Big|\big([\lambda_{t_{k}}B^{n}_{2}]\Delta E_{t_{k}}\big)\cap\sigma\Big|=$ $\Big|\left(K_{t_{k}}\Delta E_{t_{k}}\right)\cap\sigma\Big|=o(|t_{k}|),$
$\Big|\big(\left[(1+t_{k})\lambda_{t_{k}}B^{n}_{2}\right]\Delta E_{t_{k}}\big)\cap\sigma_{0}\Big|=$ $\Big|\left(K_{t_{k}}\Delta E_{t_{k}}\right)\cap\sigma_{0}\Big|=o(|t_{k}|).$
In turn, Lemma 6.2 yields that
$\Big|[\lambda_{t_{k}}B^{n}_{2}]\Delta E_{t_{k}}\Big|=o(|t_{k}|)\mbox{ \ and \ }\Big|\left[(1+t_{k})\lambda_{t_{k}}B^{n}_{2}\right]\Delta E_{t_{k}}\Big|=o(|t_{k}|),$
and hence the triangle inequality for the symmetric difference metric and $\lambda_{t_{k}}\in(\frac{3}{4},\frac{4}{3})$ imply that $\big|B^{n}_{2}\Delta\left[(1+t_{k})B^{n}_{2}\right]\Big|=o(|t_{k}|)$. However, $\big|B^{n}_{2}\Delta\left[(1+t_{k})B^{n}_{2}\right]\Big|\geq|t_{k}|\cdot\omega_{n}$, and this contradiction finally verifies (iii).
Acknowledgement: Böröczky’s research is supported in part by NKFIH NKKP grant 150613.